### v1

# Inference for concave distribution functions under measurement error

Abstract
We propose an estimator of a concave cumulative distribution function under the measurement error model, where the non-negative variables of interest are perturbed by additive independent random noise. The estimator is defined as the least concave majorant on the positive half-line of the deconvolution estimator of the distribution function. We show its uniform consistency and its square root convergence in law in $\ell_{\infty}(\mathbb{R})$. To assess the validity of the concavity assumption, we construct a test for the nonparametric null hypothesis that the distribution function is concave on the positive half-line, against the alternative that it is not. We calibrate the test using bootstrap methods. The theoretical justification for calibration led us to establish a bootstrap version of Theorem 1 in Söhl and Trabs (2012), a Donsker-type result from which we obtain, as a special case, the limiting behavior of the deconvolution estimator of the distribution function in a bootstrap setting with measurement error. Combining this Donsker-type theorem with the functional delta method, we show that the test statistic and its bootstrap version have the same limiting distribution under the null hypothesis, whereas under the alternative the bootstrap statistic is stochastically smaller. Consequently, the power of the test tends to one, for any fixed alternative, as the sample size tends to infinity. In addition to the theoretical results for the estimator and the test, we investigate their finite-sample performance in simulation studies.
Keywords. Bootstrap, Constrained estimation, Deconvolution, Functional delta method, Least Concave Majorant, Nonparametric test.
1 Introduction
Statistical inference under shape restrictions exploits nonparametric assumptions on the shape of the distribution of interest to improve estimates. Additionally, it deals with nonparametric tests about such shape constraints. The books [25, 14, 29] provide an overview of the issues addressed in the context of statistical inference under shape constraints. The results depend on the methodology and on the type of assumption, see for instance [13, 21, 8, 26, 16, 18, 17]. One of the most commonly used constraint, starting from the seminal work of [13], is the concavity of the cumulative distribution function [5, 3], corresponding to a decreasing density [14, Chapter 2.2].
Differently from the usual approach, in this paper we deal with this problem under the classic measurement error model (also known as deconvolution model), namely, we assume that the random variable of interest $X$ is contaminated by some additive random and independent noise $\varepsilon$, with a known distribution. In other words, instead of $X$, which we assume to be non-negative and absolutely continuous, we observe $Y=X+\varepsilon$. The usual way of dealing with this model is using the deconvolution estimator [31, 22]. In this paper, we propose to improve this estimator using the concavity asumption. More importantly, we will focus on testing the concavity of the cumulative distribution function of interest under the measurement error model. If we are able to confirm the concavity assumption, based on observed data, then we can use this information to improve the estimation by exploiting the shape constraint. Otherwise, if the test detects deviations from concavity, then the unconstrained estimator will be preferable. Another reason why a concavity test is particularly relevant, in the measurement error model, is that the distribution of the additive error has a crucial impact on the shape of the cumulative distribution function of $Y$. On the one hand, $X$ may have a concave distribution function but $Y$ may not (as a trivial example, among many others, if $X$ and $\varepsilon$ are uniform, the resulting $Y$ has a triangular distribution). On the other hand, in some cases when the distribution function of $X$ is non-concave, the convolution with $\varepsilon$ may yield a concave distribution function on $\mathbb{R}^{+}$.
This paper combines theories on deconvolution estimation (i.e., estimation under measurement error) and on estimation and testing under shape constraints. As a shape-constrained estimator, we consider the least concave majorant of the classic kernel deconvolution estimator of the distribution function $F$ of $X$. We show its uniform consistency under general assumptions, and we establish a Donsker-type result for it, i.e. we show that the difference between the estimator and the true concave distribution function converges in law at the rate $\sqrt{n}$ to a centered Borel random variable in $\ell_{\infty}(\mathbb{R})$. Then, inspired by [10, 8], we define our test statistic as the supremum distance between the two estimators. Although the construction is similar, the proof of the theoretical properties is much more involded in the model with measurement error than in those considered in [10, 8]. In particular, we do not have access to a least favorable hypothesis (see Section 5.2) so the calibration of the test requires challenging efforts. The performance of the constrained estimator and that of the test are illustrated through a simulation study. he simulations suggest that, for sufficiently large sample sizes and a moderate noise-to-signal ratio, (i) when $F$ is concave, the constrained estimator outperforms the unconstrained one; (ii) under concavity of $F$, the rejection rate of the test does not exceed the nominal level; and (iii) the test can detect nonconcavity of $F$ even in settings where the distribution of $Y$ is concave.
As by-products of our proofs, we provide
  1. 1.
    a bootstrap version of the Donsker Theorem 1 in [30] from which, as a particular case, one can derive the limit behavior of the deconvolution estimator of the distribution function in a bootstrap setting, under measurement error, see Theorem 7 and (14) below;
  2. 2.
    a general result on convergence of least concave majorants (and their slopes) of a sequence of continuously differentiable functions whose sequence of slopes uniformly converge, see Theorem 1 below.
The paper is organized as follows. Section 2 provides the model, definitions and assumptions. The constrained estimator is studied in Section 3, which also contains new results about uniform convergence of least concave majorants of continuously differentiable functions, as well as that of their slopes. The test statistic is defined in Section 4, where its limit behavior is studied under both the null and the alternative hypotheses. A bootstrap calibration of the test is given in Section 5, which also contains an alternative calibration that provides a less powerfull test, compared to the bootstrap, but is easier to implement. The main part of the paper ends with Section 6, which is devoted to simulations. The main sections are followed by appendices that are organized as follows. The results from Sections 2 and 3 are proved in Section A. Theorem 5, that provides the limit behavior of the test statistic, is proved in B. Section C provides a generalization of the Donsker result in [30] to a bootstrap setting. The corresponding proofs are given in Section D. Section E provides all the proofs for the bootstrap calibration.
2 Context
2.1 Notations
Let $\mathbb{R}^{+}=[0,\infty)$.
The convolution operator is denoted by $\star$.
For all $x\in\mathbb{R}$, $\langle x\rangle=(1+x^{2})^{1/2}$.
For all intervals $I\subset\mathbb{R}$, $\ell_{\infty}(I)$ denotes the space of all bounded functions $\theta:I\to\mathbb{R}$ equipped with the sup-norm $\|\theta\|_{I\infty}=\sup_{t\in I}|\theta(t)|$, where the indice $I$ will be omitted when $I=\mathbb{R}^{+}$ or when there is no possible confusion; and $C^{\infty}(I)$ denotes the set of infinitely differentiable functions on $I$. For $p\in[1,\infty)$, $L^{p}(I)$ denotes the space of real-valued functions $g$ on $I$ such that $\int_{I}|g(x)|^{p}\,dx$ is finite, equipped with the $L^{p}$-norm $\|\cdot\|_{p}$ such that for all $g$ in the space, $\|g\|_{p}^{p}$ is equal to the latter finite integral.
Weak convergence is denoted by $\rightsquigarrow$ and convergence in distribution of random variables is denoted with $\overset{L}{\longrightarrow}$.
The probability density functions of $X$, $Y$ and $\varepsilon$ are denoted by $f$, $f_{Y}$ and $f_{\varepsilon}$, the corresponding cumulative distribution functions are denoted by $F$, $F_{Y}$ and $F_{\varepsilon}$ and the corresponding Fourier transforms of are denoted by $\varphi$, $\varphi_{Y}$ and $\varphi_{\varepsilon}$, respectively. Hence in particular, $\varphi_{\varepsilon}=\mathcal{F}f_{\varepsilon}$, and similarly for the other two distributions.
We write $m\ll n$ to mean that $m$ depends on $n$ and is such that $mn^{-1}$ converges to zero as $n\to\infty$.
2.2 Model, objective and basic estimators
Under measurement error, we observe a random sample $Y_{1},\dots,Y_{n}$ from a random variable $Y=X+\varepsilon$, where $X$ is a non-negative random variable of interest and $\varepsilon$ is a random noise, which we assume to be independent from $X$. In particular, the sample can be written in the form $Y_{i}=X_{i}+\varepsilon_{i}$, for $i=1,\dots,n,$ where $X_{i}$ and $\varepsilon_{i}$ are iid copies of $X$ and $\varepsilon$, respectively.
The distribution of $X$ is not identifiable if that of $\varepsilon$ is unknown. Consequently, assuming a form for the distribution of $\varepsilon$ is frequent [11, 31, 1]. We assume that the random noise $\varepsilon$ have a known probability density function, denoted with $f_{\varepsilon}$. Moreover, we assume that $X$ has a probability density function as well. We denote the cumulative distribution function and the probability density function of $X$ with $F$ and $f$, respectively. Then, $Y$ has a probability density function that we denote by $f_{Y}$, and that is the convolution between $f$ and $f_{\varepsilon}$: for all $t\in\mathbb{R}$,
$f_{Y}(t)=f\star f_{\epsilon}(t)=\int f(t-x)f_{\varepsilon}(x)dx.$
Our aim is to build a test for the null hypothesis
$\mathcal{H}_{0}:\ ``F\text{ is concave on }[0,\infty)$", (1)
against the alternative hypothesis $\mathcal{H}_{1}:\ ``F\text{ is not concave on }[0,\infty)$”. Moreover, if the null hypothesis is accepted, we wish to construct an estimator of $F$ that takes it into account, i.e. a concave estimator. Defining the concave estimator and the test requires to first define Fourier transforms, which we will now do.
For an absolutely integrable function $g$, we define its corresponding Fourier transform as $\mathcal{F}g(t)=\int e^{its}g(s)\,ds$ for all $t\in\mathbb{R}$; note that if $Z$ is a random variable with density function $g$, then $\mathcal{F}g(t)$ represents the characteristic function $E(\exp(itZ))$ evaluated at $t$. We will estimate the density function $f$ using the so-called deconvolution approach (see, e.g., [31]), and we obtain an estimator of $F$ by integration. The deconvolution estimator is constructed by observing that, under the deconvolution model, we have $\mathcal{F}f_{Y}=\mathcal{F}f\times\mathcal{F}f_{\varepsilon}$, so, if $\mathcal{F}f_{\varepsilon}$ has no vanishing points on the real line (that is, $\mathcal{F}f_{\varepsilon}(t)\neq 0$ for all $t\in\mathbb{R}$), then $\mathcal{F}f=\mathcal{F}f_{Y}/\mathcal{F}f_{\varepsilon}$, and by applying the inverse Fourier transform $\mathcal{F}^{-1}$, we obtain
$f(x)=\mathcal{F}^{-1}\left[\frac{\mathcal{F}f_{Y}}{\mathcal{F}f_{\varepsilon}}\right](x)$
for all $x\in\mathbb{R}$. An empirical version of $\mathcal{F}f_{Y}$ can be obtained using the plug-in method combined with the kernel method, i.e., $\varphi_{n}\times\mathcal{F}K_{h_{n}}$, where $\varphi_{n}(t):=n^{-1}\sum_{i=1}^{n}\exp(itY_{i})$ and $K_{h_{n}}(x)=h_{n}^{-1}K\left(x/h_{n}\right)$, with $K$ being a kernel and $h_{n}>0$ representing the bandwidth. Consequently, the deconvolution estimator of $f$ is expressed as follows:
$f_{n}^{d}\left(x\right):=\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{n}}\frac{\varphi_{n}}{\mathcal{F}f_{\varepsilon}}\right]\left(x\right)$ (2)
for all $x\in\mathbb{R}$. Note that $f_{n}^{d}$ is finite and well defined under the usual assumptions that $\mathcal{F}K$ is compactly supported and $\mathcal{F}f_{\varepsilon}$ has no vanishing points on the real line. By integration, we get the basic estimator of $F$ as follow:
$F_{n}^{d}\left(x\right):=\int_{0}^{x}\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{n}}\frac{\varphi_{n}}{\mathcal{F}f_{\varepsilon}}\right]\left(t\right)dt$ (3)
for all $x\geq 0$. We say that it is a basic estimator because it does not satisfy a concavity constraint (concave estimators will be considered in the Section 3). By definition, the estimator takes value zero at point zero, however, it does not necessarily tend to one as a distribution function should. Hence, in some cases, we will prefer to use a normalized version of $F_{n}^{d}$, defined by
$\hat{F}_{n}^{d}(t)=\frac{F_{n}^{d}(t)}{\lim_{t\to\infty}F_{n}^{d}(t)},$ (4)
where the denominator is well defined and positive thanks to Lemma 1 below. Note that the assumptions of the lemma (which implicitely assume that $\varphi_{\epsilon}$ has no vanishing point) are satisfied under our assumptions K, D and E, but are much more general.
Lemma 1.
Let $f_{n}^{d}$ and $F_{n}^{d}$ be defined by (2) and (3). Assume that $h_{n}^{2+2\beta}n\rightarrow\infty$, $\mathcal{F}K$ is supported on [1,1][-1,1], there exist positive $\beta$ and $C$ such that $|1/\varphi_{\varepsilon}(t)|\leq C(1+|t|^{2})^{\beta/2}$ for all $t\in\mathbb{R}$, $f$ is bounded and $\int|\mathcal{F}f|$ is finite. We then have
  1. 1.
    $\|f_{n}^{d}-f\|_{\infty}=o_{p}(1)$ and $\|F_{n}^{d}-F\|_{\infty}=o_{p}(1)$.
  2. 2.
    The denominator in (4) tends to one in probability; hence it is well defined, finite and positive with probability that tends to one as $n\to\infty$.
2.3 Assumptions
A key ingredient in our proofs is Theorem 1 in Söhl and Trabs (2012). This is a Donsker-type result that gives, as a special case, the limiting behavior of the deconvolution estimator of the distribution function in the measurement error model. Assumptions K, D and E below are taken from that theorem.
Concerning the kernel $K$, we will need the following assumptions. Note that the deconvolution estimator $f_{n}^{d}$ takes real values if $K$ is symmetric and real valued, two assumptions that we assume to hold. The assumptions below are satisfied for instance by the kernel $K$ defined by (13) below where $r\geq 2$ is an even integer and $s\geq 1$ is an integer, see the third claim in Lemma 6 below.
Assumptions K
  1. 1.
    $K\in L^{1}(\mathbb{R})\cap\ell_{\infty}(\mathbb{R})$ is symmetric and band-limited with support of $\mathcal{F}K$ contained in $\left[-1,1\right].$
  2. 2.
    $\int K=1$, $\int x^{j}K\left(x\right)dx=0$ and $\int|x^{L+1}K\left(x\right)|dx<\infty,$ for j=1,,Lj=1,...,L and some integer $L>0$.
  3. 3.
    $K$ has a continuous derivative and, for a positive constant $C$, it satisfies $|K(x)|+|K^{\prime}(x)|\leq C(1+x^{2})^{-1}.$
On the other hand, we will need assumptions on the distribution of $X$. The probability density function $f$ of $X$ is assumed to satisfy the following assumptions:
Assumptions D
  1. 1.
    $f$ is bounded and satisfies $\int\left|x\right|^{2+\delta}f\left(x\right)dx<\infty$ for some positive $\delta$.
  2. 2.
    $f\in H^{\alpha}\left(\mathbb{R}\right)$, the Sobolovev space of order $\alpha\geq 0$ defined by
    $H^{\alpha}\left(\mathbb{R}\right)=\left\{f\in L_{2}(\mathbb{R})\ s.t.\ \int\left(1+x^{2}\right)^{\alpha}|\mathcal{F}f(x)|^{2}dx<\infty\right\}.$
Finally, we make assumptions on the distribution of the error. The decay rate of the characteristic function of the error $\varepsilon$} reflects the smoothness of the error. Here, we consider the case of ordinary smooth error, which means that the characteristic function of $\varepsilon$ decays with polynomial rate. The conditions are similar to the classical decay condition introduced by [11] and are satisfied by many ordinary smooth error distributions, such as the gamma or the Laplace. Specifically, we assume that:
Assumptions E
  1. 1.
    $\int\left|x\right|^{2+\delta}f_{\varepsilon}\left(x\right)dx<\infty$ for some positive $\delta$.
  2. 2.
    $\varphi_{\varepsilon}$ has no vanishing points on the real line, and for some $\beta>0$ and $C_{1}>0$, the following condition holds:
    $\left|\left(1/\varphi_{\varepsilon}\right)^{\prime}(t)\right|\leq C_{1}(1+|t|^{2})^{(\beta-1)/2},\quad\text{for all }t\in\mathbb{R}.$ (5)
3 Concave estimators
3.1 Definition of the concave estimator and consistency
A natural way of estimating $F$, exploiting the concavity assumption, is using the least concave majorant, defined below, of a basic estimator.
Definition 1.
Given an interval $I\subseteq\mathbb{R}^{+}$, the operator $\mathcal{M}_{I}:\ell^{\infty}\left(I\right)\rightarrow\ell^{\infty}\left(I\right)$ maps each function $\theta\in\ell^{\infty}\left(I\right)$ to $\mathcal{M}_{I}\theta$, the least concave majorant of $\theta$ on $I$, defined for $x\in I$ by
[…]
For convenience, we write $\mathcal{M}$ instead of $\mathcal{M}_{\mathbb{R}^{+}}$ and we call this operator “least concave majorant” without specifying the interval on which it is taken.
The following theorem, which is proved in Section A, provides a general result on convergence of least concave majorants (and their slopes) of a sequence of continuously differentiable functions whose sequence of slopes uniformly converge.
Theorem 1.
Let $f$ and $f_{n}$, for positive integers $n$, be continuous functions from $0,\infty)$ to $\mathbb{R}$ and let $F_{n}(t)=\int_{0}^{t}f_{n}(u)du$ and $F(t)=\int_{0}^{t}f(u)du$ for $t\geq 0$. Assume $F_{n},F\in\ell_{\infty}(\mathbb{R}^{+})$ and denote by $f_{n0}$ and $f_{0}$ the respective slopes of $\mathcal{M}F_{n}$ and $\mathcal{M}F$. Then one has:
  1. 1.
    $f_{n0}$ and $f_{0}$ are continuous and non-negative.
  2. 2.
    If moreover $f$ is non-negative and $\|f_{n}-f\|_{\infty}=o(1)$ as $n\to\infty$ then $\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}=o(1)$ and $\|f_{n0}-f_{0}\|_{\infty}=o(1)$.
  3. 3.
    In a stochastic setting where $f_{n}$ and $f$ are possibly random, if $f$ is non-negative and $\|f_{n}-f\|_{\infty}=o_{p}(1)$ as $n\to\infty$, then $\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}=o_{p}(1)$ and $\|f_{n0}-f_{0}\|_{\infty}=o_{p}(1)$.
As a concave estimator of $F$, we will consider the least concave majorant $\mathcal{M}F_{n}^{d}$ with $F_{n}^{d}$ denoting the basic estimator taken from (3). In some cases, we will prefer to use the least concave majorant of the normalized version $\hat{F}_{n}^{d}$ taken from (4). In particular, $\hat{F}_{n}^{d}$ seems to be more convienent for estimating $F$, while we will use $F_{n}^{d}$ for the test. Note that both estimators $\mathcal{M}F_{n}^{d}$ and $\mathcal{M}\hat{F}_{n}^{d}$ are continuous by concavity and that according to (3) and Definition 1 they take the constant value 0 on the negative half-real line. Moreover, it follows from usual properties of least concave majorants (see for instance Section 2 in [9]) that the two estimators are connected through
$\mathcal{M}\hat{F}_{n}^{d}(t)=\frac{\mathcal{M}F_{n}^{d}(t)}{\lim_{t\to\infty}F_{n}^{d}(t)}.$ (6)
As a corollary (also proved in Section A) of Theorem 1, we derive uniform convergence of our concave estimators.
Corollary 1.
Let $f_{n}^{d}$ and $F_{n}^{d}$ be defined by (2) and (3) in the deconvolution model of Section $2.$ Assume that $h_{n}^{2+2\beta}n\rightarrow\infty$, $\mathcal{F}K$ is supported on [1,1][-1,1], there exist positive $\beta$ and $C$ such that $|1/\varphi_{\varepsilon}(t)|\leq C(1+|t|^{2})^{\beta/2}$ for all $t\in\mathbb{R}$, $f$ is bounded continuous and $\int|\mathcal{F}f|$ is finite. Then
$\|\mathcal{M}_{n}-\mathcal{M}F\|_{\infty}=o_{p}(1)\quad and\quad\|f_{n0}-f_{0}\|_{\infty}=o_{p}(1)$
where $f_{0}$ is the slope of $\mathcal{M}F$, $\mathcal{M}_{n}$ is either $\mathcal{M}F_{n}^{d}$ or $\mathcal{M}\hat{F}_{n}^{d}$, and $f_{n0}$ is the slope of $\mathcal{M}_{n}$.
3.2 Limit distribution
Next, we are interested in computing the limit distribution. A key ingredient for that is Theorem 1 of [30], that provides a uniform central limit theorem for kernel estimators in the deconvolution model. The theorem covers translation classes of linear functionals of the density of $X$ of the form $t\mapsto\vartheta_{t}:=\int\zeta(x-t)f(x)dx$, where the special case $\zeta:=\mathds{1}_{(-\infty,0]}$ leads to the estimation of the distribution function $\vartheta_{t}=F(t)$ with an estimator that ressembles that in (3). Precisely, the estimator here is
$\tilde{F}_{n}^{d}\left(x\right):=\int_{-\infty}^{x}\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{n}}\frac{\varphi_{n}}{\mathcal{F}f_{\varepsilon}}\right]\left(t\right)dt$ (7)
which is connected to the estimator in (3) by $F_{n}^{d}=\tilde{F}_{n}^{d}-\tilde{F}_{n}^{d}(0)$ on $0,\infty)$. The theorem is a Donsker type of theorem in the sense that it proves that the difference between the process $(\vartheta_{t})_{t}$ and its estimator converges in law at the rate $\sqrt{n}$ to a centered Gaussian Borel random variable in $\ell_{\infty}(\mathbb{R})$. Our assumptions K, D and E are taken from that theorem. Combining Theorem 1 of [30] with Example 1 of the same paper proves the simplified following version of the theorem, for the special case $\vartheta_{t}=F(t)$.
Theorem 2.
[30] Assume conditions D and E with $\beta<1/2$, $\alpha+3\gamma>2\beta+1$ for some $\gamma\in(\beta,1/2)$ and K with $L=\left\lfloor\alpha+\gamma\right\rfloor$. Furthermore, let $h_{n}^{2\alpha+2\gamma}n\rightarrow 0$ and $h_{n}^{\rho}n\rightarrow\infty$ for some $\rho>4\beta-4\gamma+2$. Then, we have
$\sqrt{n}(\tilde{F}_{n}^{d}-F)\rightsquigarrow\mathbb{G}_{F}\ \ in\ \ell_{\infty}(\mathbb{R}),$ (8)
as $n\to\infty$, where $\mathbb{G}_{F}$ is a centered Gaussian Borel random element in $\ell_{\infty}(\mathbb{R})$, with covariance function given by
$\Sigma_{s,t}=\int g_{t}(x)g_{s}(x)f_{Y}(x)\,dx-F(t)F(s),$
where $s,t\in\mathbb{R}$ and
$g_{t}=\mathcal{F}^{-1}\left[\frac{1}{\mathcal{F}f_{\varepsilon}(-\bullet)}\right]\star\mathds{1}_{(-\infty,t]}.$
The limit distribution of the estimator $\mathcal{M}F_{n}^{d}$ from Section 3.1 can be derived from the previous theorem combined with the functional delta method by applying the functional $\mathcal{M}$. The standard version of the functional delta method assumes the operator is Hadamard differentiable (see, e.g., Theorem 2.1 in [33]), but the approach is in fact broadly applicable under Hadamard directional differentiability, see Theorem 2.1 in [28]. Hence, a key ingredient for our purpose is Proposition 2.1 in [2], according to which $\mathcal{M}$ is Hadamard directionally differentiable with a derivative in closed form. For completeness, the proposition is recalled in Theorem 3 below, where the following notations are used. For any concave function $\theta\in\ell_{\infty}(\mathbb{R}^{+})$ and $x\geq 0$, let $I_{\theta,x}$ be the union of the singleton $\{x\}$ and all open intervals $A\subset\mathbb{R}^{+}$ such that $x\in A$ and $\theta$ is affine on $A$. Note that $I_{\theta,x}=\{x\}$ if $\theta$ is not affine in a neighborhood of $x$, whereas it is an open interval with non-empty interior otherwise. Moreover, we use the notations $\mathcal{M}$ and $\mathcal{M}_{I}$ for intervals $I$ as in Definition 1, and we denote by $C_{0}$ the collection of continuous real valued functions on $\mathbb{R}^{+}$ vanishing at infinity, equipped with the sup-norm.
Theorem 3 ([2], Proposition 2.1).
The operator $\mathcal{M}$ is Hadamar directionally differentiable at any concave function $\theta\in\ell_{\infty}(\mathbb{R}^{+})$ tangentially to $C_{0}$. Its directional derivative $\mathcal{M}_{\theta}^{\prime}:C_{0}\to\ell_{\infty}(\mathbb{R}^{+})$ is uniquely determinated as follows: for any $h\in C_{0}$, and $x\geq 0$, we have $\mathcal{M}_{\theta}^{\prime}h(x)=\mathcal{M}_{I_{\theta,x}}h(x)$.
In particular, if $\theta$ is not affine in a neighborhood of $x$, then $\mathcal{M}_{\theta}^{\prime}h(x)=h(x)$ for all $h\in C_{0}$.
We are now in position to state the main result of the section, which gives the limit distribution of the concave estimator $\mathcal{M}F_{n}^{d}$. The theorem shows that the estimator converges to the least concave majorant of $F$ at the rate $\sqrt{n}$, so, in particular, the estimator converges to $F$ if the latter is concave on its support. More precisely, the difference between the estimator and the least concave majorant of $F$ converges at the rate $\sqrt{n}$ to a centered Gaussian Borel variable in $\ell_{\infty}(\mathbb{R})$.
Theorem 4.
Let the assumptions of Theorem 2 be satisfied. Then as $n\to\infty$, one has
$\sqrt{n}(\mathcal{M}F_{n}^{d}-\mathcal{M}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}(0)\quad in\ \ell_{\infty}(\mathbb{R}).$ (9)
Using that the least concave majorant of an arbitrary continuous function $G\in\ell_{\infty}(\mathbb{R}^{+})$ coincides with $G$ at point 0, one can see that the definition of $\mathcal{M}_{F}^{\prime}$ implies that $\mathcal{M}_{F}^{\prime}G(0)=G(0)$. In particular, the right-hand process in (9) (as well as the left-hand process) takes value 0 at point 0.
4 The test statistic
To test the null hypothesis (1) against the alternative that $F$ is not concave on $\mathbb{R}^{+}$, we consider a test statistic defined as a distance between the general deconvolution estimator $F_{n}^{d}$ (that performs well under both the null and the alternative hypotheses) and an estimator of $F$ that works well under the null hypothesis only, i.e. when $F$ is concave. For the latter, we consider the least concave majorant $\mathcal{M}F_{n}^{d}$ of $F_{n}^{d}$ (see Definition 1) which, according to Corollary 1, uniformly converges to $\mathcal{M}F$, which is precisely $F$ under the null hypothesis, and differs from $F$ otherwise. Hence, the test statistic typically takes small values under the null hypothesis and large values under the alternative hypothesis, so the null hypothesis will be rejected if the test statistic is too large. The test statistic is formally defined as
$T_{n}(F_{n}^{d})$ $:=$ $\sqrt{n}\sup_{t\geq 0}\left|\mathcal{M}F_{n}^{d}(t)-F_{n}^{d}(t)\right|$
$=$ $\sqrt{n}\|\mathcal{M}F_{n}^{d}-F_{n}^{d}\|_{\infty}.$
Note that due to (6) and (17) below, one has $T_{n}(\hat{F}_{n}^{d})=T_{n}(F_{n}^{d})(1+o_{p}(1))$ so replacing $F_{n}^{d}$ by $\hat{F}_{n}^{d}$ would provide a test with similar performances. Hence, for simplicity, we consider the test based on the non-normalized estimator $F_{n}^{d}$.
In order to calibrate the test, i.e. to give a precise meaning of the words “too large” above, we need to compute the limiting distribution of the test statistic under the null hypothesis. This is the aim of the following theorem, which also gives its general behavior under the alternative hypothesis. Note that under the alternative hypothesis, we are able to consider very mild assumptions since we do not need to compute the limiting distribution of the test statistic.
Theorem 5.
  1. 1.
    Let the assumptions of Theorem 2 be satisfied. If $F$ is concave on $\mathbb{R}^{+}$, then as $n\to\infty$, one has
    $T_{n}(F_{n}^{d})\overset{L}{\longrightarrow}\sup_{t\geq 0}\left|\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}(t)-\mathbb{G}_{F}(t)\right|.$ (11)
  2. 2.
    Assume that $h_{n}^{2+2\beta}n\rightarrow\infty$, $\mathcal{F}K$ is supported on [1,1][-1,1], there exist positive $\beta$ and $C$ such that $|1/\varphi_{\varepsilon}(t)|\leq C(1+|t|^{2})^{\beta/2}$ for all $t\in\mathbb{R}$, $f$ is bounded and $\int|\mathcal{F}f|$ is finite. If $F$ is not concave on $\mathbb{R}^{+}$, then there exists $C>0$ such that $T_{n}(F_{n}^{d})\geq\sqrt{n}C$ with probability that tends to one as $n\to\infty$.
An interesting property of the test that can be seen in (11) is that the limiting distribution of our test statistic does not depend on the choice of $h_{n}$. Thus, the choice of this bandwidth can be made quite arbitrarily without altering the limit behavior of the test statistic under the null hypothesis.
5 Calibration of the test
5.1 Bootstrap calibration
We will calibrate the test using bootstrap methods. For this task, we need to draw a bootstrap version of the observations in such a way that the test statistic based on those observations mimics the behavior that the test statistic based on the original observations would have had if the null hypothesis were true. Hence, the bootstrap observations should be distributed as the sum between two independent variables, one of them having a concave distribution function that is close to that of $X$ under the null hypothesis, and the other one having the same distribution as $\varepsilon$. An initial suggestion of concave distribution function that could be used to that end, is the concave estimator $\mathcal{M}F_{n}^{d}$. However, this estimator is not necessarily a genuine distribution function on $\mathbb{R}^{+}$. We say that a function $H$ is a genuine distribution function on $\mathbb{R}^{+}$ if it is a right-continuous non-decreasing function on $\mathbb{R}^{+}$ such that $H(t)=0$ for any $t\leq 0$ and $\lim_{t\to\infty}H(t)=1$. To ensure that the latter condition is fulfiled, we consider the estimator $\mathcal{M}\hat{F}_{n}^{d}$. Lemma 2 below establishes that this defines a genuine distribution function on $\mathbb{R}^{+}$.
Lemma 2.
Let the assumptions of Theorem 2 be satisfied. Then, $\mathcal{M}\hat{F}_{n}^{d}$ is a genuine distribution function on $\mathbb{R}^{+}$ with probability that tends to one.
We turn to the construction of the bootstrap observations. Fix $m\ll n$ that tends to infinity as $n\to\infty$, and conditionally on the original sample $Y_{1},\dots,Y_{n}$, consider a sample of $m$ iid random variables $X_{1}^{*},\dots,X_{m}^{*}$ with common distribution function equal to $\mathcal{M}\hat{F}_{n}^{d}$, consider a sample $\varepsilon_{1}^{*},\dots,\varepsilon_{m}^{*}$ from the same distribution as $\varepsilon$ and define $Y_{i}^{*}=X_{i}^{*}+\varepsilon_{i}^{*}$ for all $i$. Here, the sample $(\varepsilon_{i}^{*})$ is independent of all other variables $(Y_{1},\dots,Y_{n},X_{1}^{*},\dots,X_{m}^{*})$. The reason why we choose $m\ll n$ is explained in Section E.2.
Now, compute the deconvolution estimator $F_{m}^{d*}$ based on observations $Y_{1}^{*},\dots,Y_{m}^{*}$ and $T_{m}(F_{m}^{d*})$, the bootstrap version of the test statistic where the functional $T_{m}$ maps a distribution function into the supremum distance between that distribution function and its least concave majorant on $\mathbb{R}^{+}$, times $\sqrt{m}$. Let $c^{*}_{\gamma,m}$ be the quantile of order $1-\gamma$ of $T_{m}(F_{m}^{d*})$ for some fixed prescribed level $\gamma\in(0,1)$. Note that $c^{*}_{\gamma,m}$ can be computed via Monte Carlo simulations. The critical region of the test is then
$\left\{T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right\}.$ (12)
To study the properties of the test, we need to explore the limit behavior of the bootstrap version of the test statistic, both under the null and the alternative hypotheses. This is the aim of the following theorem, which is a bootstrap version of Theorem 5. A key ingredient in the proof of Theorem 5 is Theorem 2. Likewise, the proof of Theorem 6 relies on a bootstrap analogue of Theorem 2, namely Theorem 7, which is stated in Section C and proved in Section D. To prove that Assumptions B of that theorem holds in our setting, we rely on results from the broader deconvolution literature, that requires stronger assumptions than for Theorem 5. Specifically, we assume the following conditions D’ and E’ instead of D and E.
Assumptions D’
  1. 1.
    $f$ is bounded, continuous and satisfies $\int\left|x\right|^{8}f\left(x\right)dx<\infty$.
  2. 2.
    $f\in H^{\alpha}\left(\mathbb{R}\right)$ for some $\alpha>1/2$ and $\left|\left(\mathcal{F}f\right)^{(j)}(t)\right|\leq C(1+|t|)^{-j}$ for all $0\leq j\leq 4,\ t\in\mathbb{R}$.
Assumptions E’
  1. 1.
    $f_{\varepsilon}$ is bounded and satisfies $\int\left|x\right|^{8}f_{\varepsilon}\left(x\right)dx<\infty$.
  2. 2.
    $\varphi_{\varepsilon}$ has no vanishing points on the real line, and, for some $\beta>0$ and $C>0$, the condition (5) holds for all $t\in\mathbb{R}.$
We consider below a bootstrap version of the function $\tilde{F}_{n}^{d}$, see (7), that is involved in Theorem 2. Namely, we define
$\widetilde{F}_{m}^{d*}\left(x\right):=\int_{-\infty}^{x}\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{m}}\frac{\varphi_{m}^{*}}{\mathcal{F}f_{\varepsilon}}\right]\left(t\right)dt$
where $\varphi_{m}^{*}(u)=m^{-1}\sum_{j=1}^{m}e^{iuY_{j}^{*}}.$ Moreover, we make similar assumptions as in Theorem 2 with $n$ replaced by $m$. However, in order to use results from [15] on convergence of estimators of moments in the deconvolution model (see Section E.4.2), we restrict attention to the specific kernel considered in that paper. Hence, we now use a kernel $K$ such that its Fourier transform is defined by
$\mathcal{F}K(t)=(1-t^{r})^{s}\cdot\mathds{1}_{\{|t|\leq 1\}},$ (13)
where $r\geq 2$ is an even integer and $s\geq 1$ is an integer. In fact, we will assume that $r>4$ in order to ensure that $\hat{f}_{n}$ has a bounded absolute moment of order $2+\delta$, see Section E.4.4. Moreover, in order to use Corollary 1, we assume that $h_{n}^{2+2\beta}n\rightarrow\infty$ as $n\to\infty$. It follows from Lemma 6 below that the assumptions of Theorem 2 are satisfied under the assumptions of the following bootstrap version of it.
Theorem 6.
Suppose that $m$ tends to infinity as $n\to\infty$ with $m\ll n$. Assume conditions D’ and E’ with $\beta<1/2$, $\alpha+3\gamma>2\beta+1$ for some $\gamma\in(\beta,1/2)$. Assume that $K$ is given by (13) with $r>4$ and $rs>\lfloor\alpha+\gamma\rfloor$. Furthermore, let $h_{m}^{2\alpha+2\gamma}m\rightarrow 0$, $h_{m}^{\rho}m\rightarrow\infty$ for some $\rho>4\beta-4\gamma+2$, $h_{n}^{2\alpha+2\gamma}n\to 0$ and $h_{n}^{2+2\beta}n\rightarrow\infty$ as $n\to\infty$. Then, we have:
  1. 1.
    Conditionally on $(Y_{n})_{n\in\mathbb{N}}$ one has
    $\sqrt{m}(\widetilde{F}_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d})\rightsquigarrow\mathbb{G}_{F_{0}}\ \ in\ \ell_{\infty}(\mathbb{R}),$ (14)
    in probability as $n\to\infty$, where $F_{0}=\mathcal{M}F$ and $\mathbb{G}_{F_{0}}$ is as in Theorem 2 with $F_{0}$ instead of $F$.
  2. 2.
    If $F$ is concave on $\mathbb{R}^{+}$ then conditionally on $(Y_{n})_{n\in\mathbb{N}}$ one has
    $T_{m}(F_{m}^{d*})\stackrel{{\scriptstyle L}}{{\longrightarrow}}\sup_{t\geq 0}|\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}(t)-\mathbb{G}_{F}(t)|$
    in probability as $n\to\infty$.
  3. 3.
    Whether or not $F$ is concave on $\mathbb{R}^{+}$, the bootstrap test statistic $T_{m}(F_{m}^{d*})$ is bounded in probability conditionally on $(Y_{n})_{n\in\mathbb{N}}$.
To understand the meaning of the conditional weak convergence in probability in (14), one can consider the Prohorov distance, that metrizes weak convergence of probability measures on separable spaces, see e.g. [4, page 72]. The Prohorov distance $\pi(\mu,\nu)$ between two probability measures $\mu$ and $\nu$ on $\ell^{\infty}(\mathbb{R}^{+})$ (the separable space that we consider in Theorem 6 above), is defined as the infimum of those $\epsilon>0$ for which the two inequalities
$\mu(A)\leq\nu(A^{\epsilon})+\epsilon,\ \nu(A)\leq\mu(A^{\epsilon})+\epsilon$
hold for all Borel sets $A$, where
$A^{\epsilon}=\left\{f\in\ell^{\infty}(\mathbb{R}^{+}),\ \inf_{g\in A}\|f-g\|_{\infty}<\epsilon\right\}.$
This distance metrizes weak convergence in the sense that a sequence of probability measures $(\mu_{n})_{n\in\mathbb{N}}$ weakly converges to a probability measure $\mu$ on $\ell^{\infty}(\mathbb{R}^{+})$ if and only if $\pi(\mu_{n},\mu)\to 0$ as $n\to\infty$. If we denote by $\mu_{n}^{*}$ the conditionnal distribution of $\sqrt{m}(F_{m}^{d*}-\hat{F}_{n}^{d})$ and by $\mu$ the distribution of $\mathbb{G_{F}}$, then the meaning of (14) is that
$\pi(\mu_{n}^{*},\mu)\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}0,\ as\ n\to\infty.$ (15)
We now consider consistency of the test under both the null and the alternative hypotheses. Hence, we assume below that the assumptions of Theorem 6 are satisfied. Thanks to Lemma 6 below, we know that, in that case, the assumptions of Theorems 2 and 5 are satisfied as well.
Assume that $F$ is concave on $\mathbb{R}^{+}$. Then, it follows from Theorem 2 and Theorem 6 that the conditional limit distribution of $T_{m}(F_{m}^{d*})$ is the same as the limit distribution of $T_{n}(F_{n}^{d})$, so the test defined by the critical region in (12) satisfies the significance level constraint
$\lim_{n\to\infty}\mathbb{P}\left(T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right)=\gamma,$ (16)
provided that this limit distribution is continuous, see Section 23.1 in [32]. We will not attempt to rigourously prove that the limit distribution is continuous, insofar as this would constitute a separate research topic, but we give below some heuristics that are supported by our simulations below. In the case of weak concavity, where $F$ have at least an affine part, we conjecture that the limit distibution is continuous whence (16) holds. This is confirmed by numerical results in Section 6. On the other hand, if $F$ is strictly concave and nowhere affine, the limit of the test statistics is zero, by definition of $\mathcal{M}_{F}^{\prime}$. In particular, the limit distribution is not continuous. However, intuitively, the limit of the distance between an estimator and its least concave majorant tends to be stochastically larger if the underlying distribution function has affine parts (because, for an arbitrary concave function $F$, the value of the limiting distribution becomes larger if $F$ is replaced by an affine segment on a given interval where it is stricty concave). This is expected to occur more frequently in the boostrap case, where the underlying distribution function is $\mathcal{M}\hat{F}_{n}^{d}$, than in the original sample where the underlying distribution function $F$ is strictly concave. If these intuitions are correct, we may expect that the probability of rejection tends to 0, whence the test has asymptotic prescribed level $\gamma$, and is conservative in case of strict concavity. This is also confirmed by our simulation analysis in Section 6.
Assume now that $F$ is not concave. For arbitrary $C>0$ one has
$\mathbb{P}\left(T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right)$ $\geq$ $\mathbb{P}\left(c_{\gamma,m}^{*}\leq\sqrt{n}C\right)-\mathbb{P}\left(T_{n}(F_{n}^{d})\leq\sqrt{n}C\right).$
Since (39) below holds, we can apply Theorem 5, so one can choose $C>0$ such that the second probability on the right hand side tends to zero as $n\to\infty$. On the other hand, it follows from Theorem 6 that $T_{m}(F_{m}^{d*})$ is bounded in probability conditionally, which implies that $c_{\gamma,m}^{*}$ is bounded in probability, conditionally and also unconditionally. Hence, the first term on the right hand side converges to one as $n\to\infty$,
$\lim_{n\to\infty}\mathbb{P}\left(T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right)=1.$
In other words, the test defined by the critical region in (12) is consistent in the sense that it has prescribed asymptotic level, and its asymptotic power is equal to one at any point of the alternative.
5.2 About least favorable hypothesis
A common way to calibrate tests without bootstrap is to exhibit a least favorable null distribution $F_{0}$, i.e., a function satisfying the null hypothesis such that, under the null, $T_{n}(F_{n}^{d})$ is asymptotically stochastically smaller than $T_{n0}$ where $T_{n0}$ denotes the statistic obtained when the true underlying function equals $F_{0}$. Since $F_{0}$ is known, one can easily simulate $T_{n0}$ and compute its $1-\gamma$ quantile using Monte Carlo. Then, the test that rejects the null hypothesis if $T_{n}(F_{n}^{d})$ exceeds this quantile has (asymptotic) level $\gamma$, but it has smaller power than that of the test calibrated using bootstrap. Anyway, this kind of calibration is appealing since it is easier to implement than the bootstrap, so it would be interesting to see whether or not it could be implemented in our case. Unfortunately, we are not able to exibit a least favorable hypothesis in our context, and we explain why below.
Let us start with a few literature. Testing concavity (or convexity) of a cumulated hazard rate on a given bounded interval under possible censoring is considered in [8]. Testing monotonicity of a regression function, which is equivalent to testing concavity or convexity of its primitive, on a given bounded interval is considered in [10]. In both cases, the test statistic is built using least concave majorants, and affine functions are least favorable for concavity (or convexity). This is proved using the facts that the distribution of the considered estimator in [8] after centering does not depend on the function $F$ on which the test is performed, and that in [10] has the same distribution as $G_{n}\circ F$ where $G_{n}$ is a process whose distribution does not depend on $F$. Hence in both cases the way the distribution of the estimator depends on $F$ is quite simple.
Based on the literature, it is natural to wonder if affine functions are least favorable in our context. Unfortunately, it seems that it is not the case.
Indeed, all uniform distributions are affine on a bounded support, which makes them all good candidates for being least favourable. However, uniform distributions just differ in terms of scale parameters. In the classic framework, where there is no measurement error, scaling does not make any difference, so one can conveniently choose the uniform on $[0,1]$. However, in our setting this is not possible. If we rescale $X$, this corresponds to rescaling $\epsilon$, that is, changing the noise-to-signal ratio, which has a crucial impact on the estimation accuracy, and consequently, on the power of the test. Since the noise-to-signal ratio is generally unknown, it is not possible to know which uniform distribution is the least favourable in a given case. Moreover, even assuming that the variance of $X$ is known, we are not able to prove that affine functions are least favorable because the way the distribution of our estimator depends on our function $F$ is very intricate. Indeed, the limit covariance $\Sigma_{s,t}$ depends on $f_{Y}=f_{\varepsilon}\star f$ which cannot be expressed in terms of $f_{\varepsilon}\star f_{0}$, with $f_{0}$ indicating the density of the uniform distribution on the support of $f$.
A way to circumvent the absence of a least favorable hypothesis, at the cost of a slight loss of power, is to reject the null hypothesis if $T_{n}(F_{n}^{d})$ exceeds $v_{n}$, an arbitrary sequence of positive numbers that tends to infinity as $n\to\infty$ with $\sqrt{n}v_{n}$ that tends to zero. It follows from theorem 5 that $T_{n}(F_{n}^{d})$ is bounded in probability under the null hypothesis while it diverges to infinity at the rate $\sqrt{n}$ under the alternative, whence the above test has asymptotic level equal to zero and asymptotic power equal to one at any point in the alternative. The smaller $v_{n}$ is, the larger the power. One can choose for instance $v_{n}=\log n$.
6 Simulations
In this section, we implement the estimator, evaluate its performance, and assess the test by numerically calculating its power for different target distributions. To construct $f^{d}_{n}$, we employ a kernel with the following Fourier transform
$\mathcal{F}K(t)=(1-t^{6})\,\mathds{1}_{[-1,1]}(t),$
which, according to Lemma 6 , satisfies Assumption K with $L=5$. The bandwidth $h_{n}$ is chosen via the plug-in method of [7], which estimates an asymptotic approximation of the mean integrated square error to obtain a data-driven optimal bandwidth. The remainder of this section is organized as follows. In Subsection 6.1, we evaluate the performance of the proposed estimators from Section 3. In Subsection 6.2 we compute and examine the rejection rates of our test from Section 5 under different scenarios. All computations are implemented in R; the code is available at https://github.com/MohammedEssalih/Concave-CDF-under-measurement-error.
6.1 Performance of $\mathcal{M}\hat{F}_{n}^{d}$ in $\mathcal{H}_{0}$
This section evaluates the performance of the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ in terms of its mean squared error (MSE). The objective is to determine whether the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ provides improvements over its unconstrained counterpart $\hat{F}_{n}^{d}$, when $F$ is concave. To this end, we analyze the ratio of the MSE of $\mathcal{M}\hat{F}_{n}^{d}$ and $\hat{F}_{n}^{d}$ computed at different quantiles, based on $500$ simulation runs. Here, $\mathcal{M}\hat{F}_{n}^{d}$ outperforms $\hat{F}_{n}^{d}$ iif the MSE ratio is smaller than one.
Denote with $W(a,b)$ a Weibull distribution with shape and scale parameters $a>0$ and $b>0$, respectively. The distribution function of the $W(a,b)$ is concave when $a\leq 1$ and not concave for $a>1$ (for $a>1$ the function is convex to the left of an inflection point and concave to the right). We also consider the Beta distribution with shape parameters $a,b>0$, denoted with $B(a,b).$ Setting $b=1$ for simplicity, the distribution function of $B(a,1)$ is concave for $a\leq 1$, convex for $a\geq 1$, and the case $a=1$ corresponds to the uniform (the distribution function is the identity on $[0,1]$).
To ensure a comprehensive evaluation, the estimation is performed over quantile levels within the interval $(0,1)$. The bounds 0 and $1$ are omitted because the MSE of both estimators are equal to zero at those points (by definition of $\hat{F}_{n}^{d}$ and Lemma 2), whence the ratio is undefined there.
For simplicity, we restrict our attention to the case where the measurement error follows a Laplace distribution, with mean 0 and standard deviation $\sigma_{\varepsilon}$. The standard deviation is chosen to achieve a specific noise-to-signal ratio (NSR), following [6, p. 483], defined as
$\text{NSR}=\frac{\sigma_{\varepsilon}}{\sigma_{X}},$
which quantifies the degree of error contamination. Three levels of contamination, corresponding to NSR values $0.1$, $0.2$, and $0.5$, are examined across sample sizes $n\in\{20,50,100\}$.
The obtained results for the Weibull distribution $W(0.75,1)$ and the Beta distribution $B(0.75,1)$ are summarized in Figure 1.
(a) $W(0.75,1)$.
(b) $B(0.75,1)$
Figure 1: Simulation results. Mean squared error (MSE) for the Beta $B(0.75,1)$ and Weibull $W(0.75,1)$ models, across NSR levels of $10\%$, $20\%$, and $50\%$ (from left to right) and sample sizes $n=20,50,$ and $100$.
The results presented in Figure 1 show that the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ can substantially outperform its unconstrained counterpart, especially in the right tail.
6.2 Rejection rates for $T(F_{n}^{d})$
In this section, we proceed as follows. First, we describe the error distribution used in the test. Next, we implement the test in the case where the distribution function is strongly concave and nowhere affine, using the Beta and Weibull distributions. We also consider the weakly concave case, where the distribution function contains affine segments. Finally, we present an example illustrating how the presence of error can alter the shape of the distribution function, and we apply the test to examine whether it can detect this misleading effect.
6.2.1 The distribution of the error used for the test
The assumptions of Theorem 2 require a smoothness condition for the error distribution. Such a condition must hold for some $\beta<1/2$, excluding, for instance, the Laplace distribution, which corresponds to $\beta=2$.
Differently, the symmetric-gamma distribution $\text{SG}(\beta,\theta)$, with shape parameter $\beta>~0$, scale parameter $\theta>0$, and characteristic function $(1+\theta t^{2})^{-\beta}$, $t\in\mathbb{R}$, satisfies the assumptions for $\beta\in(0,1/4)$, see Example 3 of [30]. Thereore, we assume that the error distributiuon is a mixture between Laplace and symmetric-gamma, with density
$f_{\varepsilon}(x)\;=\;p\,f_{\text{SG}}\!\left(\beta,\,\theta\right)\;+\;(1-p)\,f_{\text{Lap}}\!\left(0,\,\sigma_{\text{Lap}}\right),\quad p\in(0,1),$
where $f_{\text{SG}}$ denotes the density of a symmetric-gamma distribution, and $f_{\text{Lap}}$ denotes the density of a Laplace distribution with mean 0 and scale parameter $\sigma_{\text{Lap}}>0$. It is easy to see that this mixture model satisfies the assumptions because of the symmetric-gamma component. Moreover, the Laplace component allows altering the shape of the convolution distribution function in a more effective way compared to the symmetric-gamma alone, providing more interesting cases of study. Indeed, the distribution function of the symmetric-gamma distribution is strictly concave in $(0,\infty)$, when convolved with some $F$, the resulting distribution is typically still concave regardless of the shape of $F$.
On the one hand, the choice of the symmetric-gamma parameter is not particularly critical: if $p$ is chosen very small, the overall shape of the error distribution is mainly determined by the second component of the mixture. On the other hand, the smoothness of the resulting mixture remains comparable to that of the symmetric-gamma distribution whenever the second component in the mixture is smoother than the symmetric-gamma.
To compute the NSR, the variance of $\varepsilon$ is $\operatorname{Var}(\varepsilon)=p\cdot\beta(1+\beta)\theta^{2}+(1-p)\cdot 2\sigma_{\text{Lap}}^{2}$. Setting $\theta=0.25$, $\beta=0.24$, and $p=0.01$, $\sigma_{\varepsilon}$ just depends on $\sigma_{\text{Lap}}$. To control the NSR, we choose $\sigma_{\text{Lap}}$ so that $\sigma_{\varepsilon}\;=\;\mathrm{NSR}\times\sigma_{X},$ where $\sigma_{X}$ is the standard deviation of $X$, and $\mathrm{NSR}\in\{0.1,\,0.2,\,0.5\}$.
6.2.2 The power of the test
The test performance is evaluated using $M=500$ Monte Carlo replications, $m=\lfloor n^{0.9}\rfloor$, plug-in-selected bandwidth $h_{m}$, $B=300$ bootstrap samples, and significance level $\gamma=0.1$. First, we consider $X\sim B(a,1)$ with the shape parameter $a$ ranging from 0.6 to 1.6. The results are plotted and summarized in Figure 2, with the x-axis representing $a$. Recall that the distribution function of the $B(a,b)$ is concave for $a\leq 1$ and non-concave for $a>1$, with $B(1,1)$ corresponding to the uniform.
Figure 2 indicates that, under $\mathcal{H}_{0}$ (i.e., $a\leq 1$), the rejection rates of $T(F_{n}^{d})$ remain below the nominal level $\gamma=0.1$. In particular, they tend to 0 in the strictly concave case ($a<1$) and to $\gamma=0.1$ in the affine case ($a=1$), which supports our heuristics in the paragraph after (16). Moreover, the empirical power tends to $1$ under $\mathcal{H}_{1}$ as the sample size $n$ increases. As expected, for larger (smaller) values of the NSR, the performance of the test seems to deteriorate (improve). This is quite logical, as it becomes harder to detect which hypothesis is true when the measurement error gets larger. We also examine the Weibull distribution $W(a,b)$, which belongs to $\mathcal{H}_{0}$ for $a\leq 1$ and to $\mathcal{H}_{1}$ otherwise. Figure 2(b) shows a similar pattern, with the distinction that this cumulative distribution has no affine segments. Here, the power of the test approaches 0 for $a\leq 1$ and $1$ for $a>1$.
(a) $\mathrm{Beta}(a,1)$ with $a=0.6,0.8,\ldots,1.6$.
(b) Weibull $W(a,1)$ with $a=0.6,0.8,\ldots,1.6$.
Figure 2: Rejection rates for $\mathrm{Beta}(a,1)$ and $W(a,1)$ distributions for different values of $a$, NSR, and $n$.
We now consider a second scenario within $\mathcal{H}_{0}$, namely, a distribution function which is affine on a non-empty interval and strictly concave elsewhere. To illustrate this, we consider the following mixture $X\sim 1/2\cdot\mathrm{U}[0,1]\;+\;1/2\cdot\bigl(1+\mathrm{Exp}(1)\bigr),$ where $\mathrm{U}[0,1]$ denotes the uniform distribution on $[0,1]$ $\mathrm{U}[0,1]$ and $\mathrm{Exp}(1)$ is the standard exponential distribution with scale parameter equal to 1. This mixture distribution is affine on $[0,1]$ and strictly concave on $(1,\infty)$. The simulation results, reported in Table 1, show that the rejection rates converge to the significance level $\gamma$ as $n$ increases, which also supports our heuristics in the paragraph after (16).
$n$ NSR = 0.1 NSR = 0.2 NSR = 0.5
100 0.115 0.085 0.060
200 0.103 0.095 0.080
500 0.1005 0.0995 0.098
Table 1: Rejection rates for the mixture $X\sim 0.5\cdot\mathrm{U}[0,1]\;+\;0.5\cdot(1+\mathrm{Exp}(1))$ distribution for different sample sizes $n$ and NSR values.
6.3 The performance of the test when the error alters the shape
In some situations, the presence of measurement error affects the shape of the convolution substantially. For example, when $X$ follows the mixture distribution $X\sim 0.2\,W(3,1)+0.8\,B(0.5,0.75),$ then $F$ satisfies $\mathcal{H}_{1}$. However, after convolution with measurement error, the distribution function of $X+\varepsilon$ may fall into $\mathcal{H}_{0}$, especially when the NSR is large. In such cases, it is essential to apply the proposed test in order ton discover the shape of the true underlying distribution $F$.
The performance of the test is reported in Table 2 for different NSR levels and sample sizes $n$. The results show that the NRS has a big impact on the performance. However, as $n$ increases, the test is able to detect the non-concavity of $F$, especially for moderate NRS values. In situations like this, our test is especially useful, as it helps discovering the true shape of a distribution function, removing the impact of measurement error.
NSRNSR $n=20$ $n=50$ $n=100$ $n=200$ $n=500$ $n=1000$
$0.1$ 0.012 0.063 0.109 0.211 0.594 0.781
$0.2$ 0.018 0.051 0.082 0.147 0.358 0.601
$0.5$ 0.026 0.039 0.045 0.072 0.101 0.176
Table 2: Rejection rates for the mixture $X\sim 0.2\,W(3,1)+0.8\,B(0.5,0.75)$ distribution for different sample sizes $n$ and NSRNSR values.
Appendix A Proofs for Sections 2 and 3
A.1 Proof of Lemma 1
We write $h$ insted of $h_{n}$. The proof relies on the results from [19] concerning the strong consistency of density estimators. Since $\mathcal{F}K$ is supported on [1,1][-1,1], the estimator in (2) aligns with the truncated density estimator in [19]. Specifically, if the Fourier transform $\mathcal{F}K$ is compactly supported in $\left[-1,1\right]$, with $K_{h}(x)=h^{-1}K(x/h)$, then $\mathcal{F}K_{h}(t)=\mathcal{F}K(th)$ is also compactly supported, with support in $\left[-1/h,1/h\right]$. Consequently, the estimator in (2) can be expressed with a finite integral instead of over $(-\infty,+\infty)$:
$f_{n}^{d}\left(x\right)=\frac{1}{2\pi}\int_{-M}^{M}\exp\left(-itx\right)\mathcal{F}K\left(th\right)\frac{\varphi_{n}\left(t\right)}{\mathcal{F}f_{\varepsilon}\left(t\right)}dt\text{ where }M=1/h.$
Arguing as in the proof of Theorem 3.1 in [19] with $M=1/h$ proves that
$\mathbb{E}\int_{-M}^{M}|\mathcal{F}f_{n}^{d}(t)-\mathcal{F}f(t)|dt\leq\frac{M\sqrt{2}}{\pi\sqrt{n}\min_{|t|\leq M}|\varphi_{\varepsilon}(t)|}+o(1),$
and that
$\mathbb{E}\|f_{n}^{d}-f\|_{\infty}\leq\frac{1}{2\pi}\mathbb{E}\int_{-M}^{M}|\mathcal{F}f_{n}^{d}(t)-\mathcal{F}f(t)|dt+\int_{|t|>M}|\mathcal{F}f(t)|\,dt.$
Under our assumptions one has $1/|\varphi_{\varepsilon}(t)|\leq CM^{\beta}$ for some positive constant $C$ and all $t\in[-M,M]$ so we have
$\frac{M}{\sqrt{n}\min_{|t|\leq M}|\varphi_{\varepsilon}(t)|}\leq C\frac{M^{1+\beta}}{\sqrt{n}}$
where the right-hand side converges to zero under the assumption that $h^{2+2\beta}n\rightarrow\infty$. Moreover, $\int|\mathcal{F}f(t)|\,dt$ is finite by assumption, so $\int_{|t|>M}|\mathcal{F}f(t)|\,dt$ converges to zero as $n\to\infty$ and we get
$\mathbb{E}\|f_{n}^{d}-f\|_{\infty}=o(1).$
It then follows from Markov’s inequality that
$\|f_{n}^{d}-f\|_{\infty}=o_{p}(1).$
Moreover, for all $t\geq 0$, one has
$|F^{d}_{n}(t)-F(t)|\leq\int_{0}^{t}|f^{d}_{n}(u)-f(u)|du\leq t\|f^{d}_{n}-f\|_{\infty},$
where the supremum on the right-hand side tends to zero. Then $F^{d}_{n}(t)$ converges to $F(t)$ for every $t\in\mathbb{R}^{+}$. Furthermore, the convergence is uniform over $\mathbb{R}^{+}$ by applying similar arguments as in the proof of Lemma 2.11 in [32], using that $F$ is monotone (by non-negativity of $f$), continuous and bounded. This completes the proof of the first claim.
Now, $F_{n}^{d}$ is continuous on $\mathbb{R}^{+}$ so $\lim_{t\to\infty}F_{n}^{d}(t)$ exists in $\infty$$\infty$] and we have
$|\lim_{t\to\infty}F_{n}^{d}(t)-\lim_{t\to\infty}F(t)|$ $\leq$ $\|F_{n}^{d}-F\|_{\infty}=o_{p}(1),$
where $\lim_{t\to\infty}F(t)=1$. Therefore,
$\lim_{t\to\infty}F_{n}^{d}(t)=1+o_{p}(1),$ (17)
where the right-hand side is finite and positive with probability that tends to one as $n\to\infty$. This proves the second claim of the lemma. $\Box$
A.2 Proof of Theorem 1
We prove the first claim for $f_{n0}$ only, since the proof for $f_{0}$ is similar. If there exists $x>0$ where $f_{n0}$ is not continuous, then the left-hand slope of $\mathcal{M}F_{n}$ at $x$ is strictly larger than its right-hand slope. This implies that $\mathcal{M}F_{n}(x)=F_{n}(x)$ and because $\mathcal{M}F_{n}$ is above $F_{n}$ by definition, this implies that the left-hand slope of $F_{n}$ at $x$ is strictly larger than its right-hand slope, which contradicts differentiability of $F_{n}$. Hence, $f_{n0}$ is continuous.
We now prove that $f_{n0}$ is non-negative. Equivalently we prove that $\mathcal{M}F_{n}$ is non-decreasing. Arguing by contradiction, we assume that $\mathcal{M}F_{n}$ is not non-decreasing. Then there exists $y_{n}>0$ where the slope $u_{n}$ of $\mathcal{M}F_{n}$ is stricly negative, whence $u_{n}>0$. By concavity of $\mathcal{M}F_{n}$ we then have
$\mathcal{M}F_{n}(y_{n}+t)\leq\mathcal{M}F_{n}(y_{n})-u_{n}t$
for all $t>0$ whence $\lim_{t\to\infty}\mathcal{M}F_{n}(t)=$-$\infty$. Since $\mathcal{M}F_{n}$ is above $F_{n}$ by definition, this implies that $\lim_{t\to\infty}F_{n}(t)=$-$\infty$ which contradicts the assumption that $F_{n}$ is bounded, whence $f_{n0}$ is non-negative.
We turn to the proof of the second claim. For all $t\geq 0$, one has
$|F_{n}(t)-F(t)|\leq\int_{0}^{t}|f_{n}(u)-f(u)|du\leq t\sup_{u\in\mathbb{R}^{+}}|f_{n}(u)-f(u)|,$
where by assumption, the supremum on the right-hand side tends to zero. Then, for every $t\in\mathbb{R}^{+}$, $F_{n}(t)$ converges to $F(t)$. Furthermore, the convergence is uniform over $\mathbb{R}^{+}$ by applying similar arguments as in the proof of Lemma 2.11 in [32], using that $F$ is monotone (by non-negativity of $f$), continuous and bounded. Then, it follows from Marshall’s inequality [14, Exercise 3.1] that
$\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}\leq\|F_{n}-F\|_{\infty},$
whence the left-hand side converges to zero.
Now, we show that $f_{n0}$ uniformly converges to $f_{0}$. It follows from the max-min formulas that for arbitrary $x>0$ one has
$f_{n0}(x)=\inf_{v>x}\sup_{u<x}\left(\frac{1}{v-u}\int_{u}^{v}f_{n}(t)dt\right)\mbox{ and }f_{0}(x)=\inf_{v>x}\sup_{u<x}\left(\frac{1}{v-u}\int_{u}^{v}f(t)dt\right),$
and for all $u<v$. Morever, the triangle inequality yields that
$\left|\frac{1}{v-u}\int_{u}^{v}f_{n}(t)dt-\frac{1}{v-u}\int_{u}^{v}f(t)dt\right|$ $\leq$ $\frac{1}{v-u}\int_{u}^{v}|f_{n}(t)-f(t)|dt$
$\leq$ $\|f_{n}-f\|_{\infty}=o(1).$
Hence,
$f_{n0}(x)$ $=$ $\inf_{v>x}\sup_{u<x}\left(\frac{1}{v-u}\int_{u}^{v}f_{n}(t)dt\right)$
$=$ $\inf_{v>x}\sup_{u<x}\left(\frac{1}{v-u}\int_{u}^{v}f(t)dt\right)+o(1)$
$=$ $f_{0}(x)+o(1).$
For $x=0$ the min max formulas become
$f_{n0}(0)=\inf_{v>0}\left(\frac{1}{v}\int_{0}^{v}f_{n}(t)dt\right)\mbox{ and }f_{0}(0)=\inf_{v>0}\left(\frac{1}{v}\int_{0}^{v}f(t)dt\right),$
and similar as above, the triangle inequality implies that for all $v>0$ one has
$\left|\frac{1}{v}\int_{0}^{v}f_{n}(t)dt-\frac{1}{v}\int_{0}^{v}f(t)dt\right|$ $\leq$ $\|f_{n}-f\|_{\infty}=o(1).$
Hence,
$f_{n0}(0)$ $=$ $\inf_{v>0}\left(\frac{1}{v}\int_{0}^{v}f_{n}(t)dt\right)$
$=$ $f_{0}(0)+o(1).$
We have proved that $f_{n0}(x)$ converges to $f_{0}(x)$ for all $x\geq 0$. The uniform convergence of $f_{n0}$ follows by continuity, boundedness and monotonicity of $f_{0}$ with similar arguments as in the proof of Lemma 2.11 in [32].
Consider now the stochastic setting. If $\|f_{n}-f\|_{\infty}$ converges to zero in probability, then from every subsequence one can extract a further subsequence along which $\|f_{n}-f\|_{\infty}$ converges to zero almost surely. Along such a subsequence, we know that $\|f_{n0}-f\|_{\infty}$ and $\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}$ converge to zero almost surely. Hence from every subsequence one can extract a further subsequence along which $\|f_{n0}-f_{0}\|_{\infty}$ and $\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}$ converge to zero almost surely, whence in probability; whence the convergences hold along the whole sequence in probability. This completes the proof of Theorem 1. $\Box$
A.3 Proof of Corollary 1
It follows from Lemma 1 that $\|f_{n}^{d}-f\|_{\infty}=o_{p}(1)$ and $F_{n}^{d}\in\ell_{\infty}(\mathbb{R}^{+})$ with probability that tends to one. Hence, Theorem 1 ensures the announced results when $\mathcal{M}_{n}$ is $\mathcal{M}F_{n}^{d}$. The results when $\mathcal{M}_{n}$ is $\mathcal{M}\hat{F}_{n}^{d}$ easily follow using (17) and (6). $\Box$
A.4 Proof of Theorem 4
We have $F_{n}^{d}=\tilde{F}_{n}^{d}-\tilde{F}_{n}^{d}(0)$ by definition, and $F(0)=0$, so it follows from Theorem 2 that
$\sqrt{n}(F_{n}^{d}-F)$ $=$ $\sqrt{n}(\tilde{F}_{n}^{d}-F)-\sqrt{n}(\tilde{F}_{n}^{d}-F)(0)$
$\rightsquigarrow$ $\mathbb{G}_{F}-\mathbb{G}_{F}(0)\quad in\ \ell_{\infty}(\mathbb{R}).$
Combining with the functional delta method and Theorem 3 yields
$\sqrt{n}(\mathcal{M}F_{n}^{d}-\mathcal{M}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}(\mathbb{G}_{F}-\mathbb{G}_{F}(0))\ \ in\ \ell_{\infty}(\mathbb{R}),$
as $n\to\infty$. The result follows since for arbitrary function $G\in C_{0}$ and constant $c\in\mathbb{R}$ one has $\mathcal{M}_{F}^{\prime}(G+c)=\mathcal{M}_{F}^{\prime}G+c$. $\Box$
Appendix B Proof of Theorem 5
Define the operator $\mathcal{W}=\mathcal{M}$-$\mathcal{I}$ where $\mathcal{I}$ stands for the identity operator. Assume $F$ is concave on $\mathbb{R}^{+}$. Then $\mathcal{M}F=F$, so the statistic in (4) becomes
$T_{n}(F_{n}^{d})=\sqrt{n}\|\mathcal{W}F_{n}^{d}-\mathcal{W}F\|_{\infty}.$ (18)
Now, for an arbitrary function $\theta\in\ell_{\infty}(\mathbb{R}^{+})$ and $a\in\mathbb{R}$ one has $\mathcal{M}(\theta+a)=\mathcal{M}\theta+a$, which implies that $\mathcal{W}(\theta+a)=\mathcal{W}\theta$. Since by definition one has $F^{d}_{n}=\tilde{F}_{n}^{d}-\tilde{F}_{n}^{d}(0)$ where $\tilde{F}_{n}^{d}$ is taken from (7), one gets
$T_{n}(F_{n}^{d})=\sqrt{n}\|\mathcal{W}\tilde{F}_{n}^{d}-\mathcal{W}F\|_{\infty}.$ (19)
Now, from Theorem 3, we can deduce that $\mathcal{W}$ is also Hadamard directionally differentiable at $F$ (given that $F$ is assumed to be concave) and its derivative is $\mathcal{W}_{F}^{\prime}=\mathcal{M}_{F}^{\prime}$-$\mathcal{I}$ since the Hadamard directional derivative of the identity is itself. Hence, combining Theorem 2 with the functional delta method and Theorem 3 yields
$\sqrt{n}(\mathcal{W}\tilde{F}_{n}^{d}-\mathcal{W}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}\ \ in\ \ell_{\infty}(\mathbb{R}),$
as $n\to\infty$. Since the sup-norm is a continuous function on $\ell_{\infty}(\mathbb{R})$, invoking the continuous mapping theorem completes the proof of the first claim in Theorem 5.
Now, assume that $F$ is not concave on $\mathbb{R}^{+}$ and define
$A:=\|F-\mathcal{M}F\|_{\infty},$
whence $A$ is strictly positive. It follows from the triangle inequality that
$n^{-1/2}T_{n}(F_{n}^{d})$ $\geq$ $\|F-\mathcal{M}F\|_{\infty}-\|F-F_{n}^{d}\|_{\infty}-\|\mathcal{M}F_{n}^{d}-\mathcal{M}F\|_{\infty}.$
Using Marshall’s inequality [14, Exercise 3.1] on gets
$n^{-1/2}T_{n}(F_{n}^{d})$ $\geq$ $\|F-\mathcal{M}F\|_{\infty}-2\|F-F_{n}^{d}\|_{\infty}$
$\geq$ $A-2\|F-F_{n}^{d}\|_{\infty}.$
Therefore,
$\mathbb{P}\left(n^{-1/2}T_{n}(F_{n}^{d})\leq A/2\right)$ $\leq$ $\mathbb{P}\left(\|F-F_{n}^{d}\|_{\infty}>A/4\right).$
Now, it follows from Lemma 1 that $\|F-F_{n}^{d}\|_{\infty}$ converges in probability to zero and therefore,
$\lim_{n\to\infty}\mathbb{P}\left(\|F-F_{n}^{d}\|_{\infty}>A/4\right)=0.$
The second claim in Theorem 5 follows from the previous two displays by choosing $C=A/2$. $\Box$
Appendix C Generalization of the Donsker theorem in [30] to a bootstrap setting
Recall that in the setting of [30], one observes independent and identically distributed random variables $Y_{j}=X_{j}+\varepsilon_{j}$ for $j=1,\dots,n$ where $X_{j}$ and $\varepsilon_{j}$ are independent of each other, the distribution of $\varepsilon_{j}$ is supposed to be known and the aim is statistical inference on the distribution of $X_{j}$. In this deconvolution model, the random variables $X_{1},\dots,X_{n}$ are not observed, but only $Y_{1},\dots,Y_{n}$, which contains an additive error. In this model, kernel estimators $\widehat{\vartheta_{t}}$ are considered for linear functionals
$t\mapsto\mathcal{\vartheta}_{t}:=\int\zeta(x-t)f_{X}(x)dx,$
with $f_{X}$ being the probability density function of the $X_{i}$s, where the special case $\zeta:=\mathds{1}_{(-\infty,0]}$ leads to the estimation of the distribution function of the $X_{i}$s. Theorem 1 in [30] states that $\sqrt{n}(\widehat{\vartheta_{t}}-\vartheta_{t})_{t\in\mathbb{R}}$ converges in law to a centered Gaussian process $\mathbb{G}$ in $\ell^{\infty}(\mathbb{R})$, whose covariance depends on the functional $\zeta$ and on the distribution of the errors $\varepsilon_{j}$. Here,
$\widehat{\vartheta_{t}}=\int\zeta(x-t)\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{n}}\frac{\varphi_{n}}{\varphi_{\epsilon}}\right](x)dx,$
where $K_{h_{n}}$ and $\varphi_{n}$ are defined as in (2), and $\varphi_{\varepsilon}$ denotes the Fourier transform of the density of $\varepsilon$.
The aim of the current section is to establish a bootstrap version of that theorem. Hence, we assume that we have at hand an estimator $\hat{f}_{n}$ for the density of the $X_{j}$s based on observations $Y_{1},\dots,Y_{n}$. With $(m)$ a sequence of positive integers that may depend on $n$, conditionally on $Y_{1},\dots,Y_{n}$, we build independent and identically distributed variables $X_{1}^{*},\dots,X_{m}^{*}$ with density function $\hat{f}_{n}$. (Note that we assume here that $\hat{f}_{n}$ is a genuine density function, which means that it is non-negative and integrates to one). Independently of all those variables, we build independent and identically distributed variables $\varepsilon_{1}^{*},\dots,\varepsilon_{m}^{*}$ with the same distribution as the original errors $\varepsilon_{j}$s. Then, conditionally on the original observations $Y_{1},\dots,Y_{n}$, we obtain independent and identically distributed bootstrap observations
$Y_{j}^{*}=X_{j}^{*}+\varepsilon_{j}^{*},\quad j=1,\dots,m.$
We then define the estimator in the bootstrap world as follows
$\widehat{\vartheta_{t}^{*}}=\int\zeta(x-t)\mathcal{F}^{-1}\left[\mathcal{F}K_{h_{m}}\frac{\varphi_{m}^{*}}{\varphi_{\epsilon}}\right](x)dx,$
where $\varphi_{m}^{*}(u)=m^{-1}\sum_{j=1}^{m}e^{iuY_{j}^{*}},$ $u\in\mathbb{R}$ and $\zeta:\mathbb{R}\to\mathbb{R}$ is a given function that belongs to $Z^{\gamma_{s},\gamma_{c}}$ for some $\gamma_{s},\gamma_{c}>0$, where (as in [30]) $Z^{\gamma_{s},\gamma_{c}}$ is the set of functions $\zeta$ such that $\zeta=\zeta^{c}+\zeta^{s}$ for some $\zeta^{s}\in H^{\gamma_{s}}(\mathbb{R})$ that is compactly supported as well as $\langle x\rangle^{\tau}(\zeta^{c}(x)-a(x))\in H^{\gamma_{c}}(\mathbb{R})$ for some $\tau>0$ and some $a\in C^{\infty}(\mathbb{R})$ such that $a^{\prime}$ is compactly supported. Similar to Assumption D in Section 2.3 (which is Assumption 1 in [30]) we make the following assumptions for the bootstrap version.
Assumption B
  1. 1.
    The estimator $\hat{f}_{n}$ is a genuine density function such that
    $\sup_{x}\hat{f}_{n}(x)=O_{P}(1)$ (20)
    and
    $\int|x|^{2+\delta}\hat{f}_{n}(x)dx=O_{P}(1)$ (21)
    for some arbitrarily small $\delta>0$.
  2. 2.
    One has
    $\int\langle u\rangle^{2\alpha}|\mathcal{F}\hat{f}_{n}(u)|^{2}du=O_{p}(1)$ (22)
    for some $\alpha>0$, where $\langle u\rangle=(1+u^{2})^{1/2}$.
  3. 3.
    $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero in probability as $n\to\infty$.
Under these assumptions, we have the following Donsker theorem, which generalizes Theorem 1 in [30] to a general bootstrap setting.
Theorem 7.
Suppose that $m$ tends to infinity as $n\to\infty$. Assume conditions E and B as well as $\zeta\in Z^{\gamma_{s},\gamma_{c}}$ with $\gamma_{s}>\beta$, $\gamma_{c}>(1/2\vee\alpha)+\gamma_{s}$ and $\alpha+3\gamma_{s}>2\beta+1$. Furthermore, let the kernel $K$ satisfy conditions K with $L=\left\lfloor\alpha+\gamma_{s}\right\rfloor$. Let $h_{m}^{2\alpha+2\gamma_{s}}m\rightarrow 0$ and if $\gamma_{s}\leq\beta+1/2$ let in addition $h_{m}^{\rho}m\rightarrow\infty$ for some $\rho>4\beta-4\gamma_{s}+2$. Then, contionally on $(Y_{n})_{n\in\mathbb{N}}$ one has
$\sqrt{m}(\widehat{\vartheta_{t}^{*}}-\vartheta_{t}^{*})_{t\in\mathbb{R}}\rightsquigarrow\mathbb{G}_{F}\ \ in\ \ell_{\infty}(\mathbb{R})$
in probability as $n\to\infty$, where $\vartheta_{t}^{*}=\int\zeta(x-t)\hat{f}_{n}(x)dx$ and $\mathbb{G}_{F}$ is the same centered Gaussian Borel random element in $\ell^{\infty}(\mathbb{R})$ as in Theorem 2.
Appendix D Proofs for Section C
The proof of Theorem 7 is given in a first subsection that also states a few lemmas that are required for the proof of the theorem. The lemmas are proved in separate subsections afterwards.
D.1 Proof of Theorem 7
For notational convenience (to have a single index that tends to infinity), we consider only the case where $m=n$; the general case can be obtained likewise. Moreover, we write $h$ instead of $h_{m}$ to alleviate the notation.
Our aim is to show that $\pi(\mu_{n}^{*},\mu)$ converges in probability to zero as $n\to\infty$, where $\mu_{n}^{*}$ is the conditional distribution of $\sqrt{n}(\widehat{\vartheta}^{*}_{t}-\vartheta_{t}^{*})_{t\in\mathbb{R}}$, $\mu$ is the distribution of $\mathbb{G}_{F}$ and $\pi$ is the Prohorov distance, see the comments following Theorem 6. We will prove below that this holds if, instead of assuming that $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero in probability, we make the stronger assumption that this convergence holds almost surely. This suffices for the following reason. If $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero in probability, then from every subsequence one can extract a further subsequence along which $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero almost surely. Along that further subsequence, we know that $\pi(\mu_{n}^{*},\mu)$ converges in probability to zero. Hence from every subsequence one can extract a further subsequence along which $\pi(\mu_{n}^{*},\mu)$ converges in probability to zero; whence the convergence holds along the whole sequence.
Hence, in the sequel we assume without loss of generality that $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero almost surely. For some fixed $C>0$, we denote by $E_{n}$ the event
$E_{n}:=\left(\sup_{t}|\hat{f}_{n}(t)|\leq C\right)\cap\left(\int\langle x\rangle^{2+\delta}\hat{f}_{n}(x)dx\leq C\right)\cap\left(\int\langle u\rangle^{2\alpha}|\mathcal{F}\hat{f}_{n}(u)|^{2}du\leq C\right).$
We will show below that if $E_{n}$ holds for all $n\in\mathbb{N}$, then conditionally on the original observations,
$\sqrt{n}(\widehat{\vartheta_{t}^{*}}-\vartheta_{t}^{*})_{t\in\mathbb{R}}\rightsquigarrow\mathbb{G}_{F}\ in\ \ell^{\infty}(\mathbb{R}).$ (23)
This means that $\mu_{n}^{*}$ weakly converges to $\mu$ so that $\lim_{n\to\infty}\pi(\mu_{n}^{*},\mu)\mathds{1}_{E_{n}}=0$, which implies that $\pi(\mu_{n}^{*},\mu)\mathds{1}_{E_{n}}$ converges in probability to zero. Hence, for all $\epsilon>0$ one has
$\mathbb{P}(\pi(\mu_{n}^{*},\mu)>\epsilon)$ $\leq$ $\mathbb{P}(\pi(\mu_{n}^{*},\mu)\mathds{1}_{E_{n}}>\epsilon)+(1-\mathbb{P}(E_{n}))$
$=$ $o(1)+1-\mathbb{P}(E_{n}).$
Now, it follows from Assumption B that for arbitrary $\eta>0$, one can choose $C>0$ large enough so that $\mathbb{P}(E_{n})\geq 1-\eta$ for sufficiently large $n$ and therefore,
$\lim_{n\to\infty}\mathbb{P}(\pi(\mu_{n}^{*},\mu)>\epsilon)$ $\leq$ $\eta.$
Letting $\eta\to 0$ proves that for all $\epsilon>0$ one has
$\lim_{n\to\infty}\mathbb{P}(\pi(\mu_{n}^{*},\mu)>\epsilon)=0,$
which proves the theorem.
It remains to prove (23), assuming that $E_{n}$ holds for all $n$ and that $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero almost surely. For this task, for an arbitrary fixed $t\in\mathbb{R}$ we decompose the error into a stochastic error term and a bias term as follows
$\widehat{\vartheta_{t}^{*}}-\vartheta_{t}^{*}=S_{nt}+B_{nt}$
where
$S_{nt}=\int\zeta(x-t)\mathcal{F}^{-1}\left[\mathcal{F}K_{h}\frac{\varphi_{n}^{*}-\hat{\varphi}_{n}}{\varphi_{\epsilon}}\right](x)dx,$
and
$B_{nt}=\int\zeta(x-t)(K_{h}\star\hat{f}_{n}-\hat{f}_{n})(x)dx.$
Here, $\hat{\varphi}_{n}$ denotes the Fourier transform of the distribution of the $Y_{j}^{*}$s. We recall that the distribution of the $\varepsilon_{i}^{*}$s is the same as that of the $\varepsilon_{i}$s whence the bootstrap version of $\varphi_{\epsilon}$ is $\varphi_{\epsilon}$ itself and therefore,
$\hat{\varphi}_{n}(u)=\varphi_{\epsilon}(u)\int e^{iux}\hat{f}_{n}(x)dx.$
Since we have assumed that $E_{n}$ holds for all $n,$ we have the uniform bounds
$\sup_{t}|\hat{f}_{n}(t)|\leq C,\quad\int\langle u\rangle^{2\alpha}|\mathcal{F}\hat{f}_{n}(u)|^{2}du\leq C\quad\mbox{and}\quad\int|x|^{2}\hat{f}_{n}^{2}(x)dx\leq C,$
so we can conclude as in [30, Section 4.1.1.], where the only difference is that $f_{X}$ has to be replaced by $\hat{f}_{n}$ that is uniformly controlled, that the bias term $B_{n}$ satisfies
$\sup_{t}|B_{nt}|=o(n^{-1/2}).$
Hence, (23) is equivalent to
$(\sqrt{n}S_{nt})_{t\in\mathbb{R}}\rightsquigarrow\mathbb{G}_{F}\ in\ \ell^{\infty}(\mathbb{R}).$ (24)
To deal with the stochastic term $S_{n}$ we use the smoothed adjoint inequality (13) in [30] to obtain that (similar to Equation (15) in that paper)
$S_{nt}=\int\ \mathcal{F}^{-1}\left[\varphi_{\epsilon}^{-1}(-\ {\bullet})\mathcal{F}K_{h}\right|\star\zeta_{t}(x)(\mathbb{P}_{n}^{*}-\widehat{\mathbb{P}}_{n})(\mbox{d}x),$ (25)
where $\zeta_{t}(x)=\zeta(t-x)$ for all $x\in\mathbb{R}$, $\mathbb{P}_{n}^{*}$ is the empirical distribution of the bootstrap sample $Y_{1}^{*},\dots,Y_{n}^{*}$, and $\widehat{\mathbb{P}}_{n}$ is the common distribution of the bootstrap observations $Y_{i}^{*}$s conditionally on the original observations $Y_{j}$s, that is the distribution with density $\hat{f}_{n}\star f_{\varepsilon}$ with $f_{\varepsilon}$ being the density of the $\varepsilon_{i}$s.
Proving (24) taking advantage of the above empirical form for $S_{n}$ generalizes arguments from the proof of [30, Theorem 1] whence details are omitted and we only point out the main steps and differences (which comes from the fact that in our case, the underlying distribution of the observations depends on $n$). Denoting by $\nu_{n}^{*}$ the conditionnal distribution of $\sqrt{n}S_{n}$, we will apply [33, Theorem 1.5.4.] according to which it suffices to prove that the finite dimensional marginals of $\nu_{n}^{*}$ converge to that of $\mu$, and that the sequence $(\nu_{n}^{*})_{n\in\mathbb{N}}$ is asymptotically tight, in the sense that for all $\epsilon>0$ there exists a compact $K\subset\ell_{\infty}(\mathbb{R})$ such that
$\liminf_{n\to\infty}\nu_{n}^{*}(K^{\eta})\geq 1-\epsilon\quad\mbox{ for all }\eta>0,$
where $K^{\eta}$ is the set of all functions $f\in\ell_{\infty}(\mathbb{R})$ such that $\|f-K\|_{\infty}<\eta$.
Convergence of the finite dimensional marginals is given in the Lemma 3 below.
Lemma 3.
Under the assumptions of Theorem 7, if $E_{n}$ holds for all $n$ and some small enough $\delta>0$, and $\|\hat{f}_{n}-f\|_{\infty}$ converges to zero almost surely as $n\to\infty$, then the finite dimensional distributions of $\sqrt{n}S_{n}$ converge towards that of $\mathbb{G}_{F}$, conditionally on the original observations.
To prove asymptotic tightness, we split the process into three parts
$\sqrt{n}S_{n}=\sqrt{n}\int(T_{1}(x)+T_{2}(x)+T_{3}(x))({\mathbb{P}}_{n}^{*}-\widehat{\mathbb{P}}_{n})(\mbox{d}x)$
where we define $T_{1}$, $T_{2}$ and $T_{3}$ in the same way as in Section 4.2.2 of [30], see their Equations (28), (29) and (30). As mentionned in that paper, due to [33, Theorem 1.5.7.], it suffices to show that the three processes $\sqrt{n}\int T_{j}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ are asymptotically tight, for $j\in\{1,2,3\}$.
In Section 4.2.2 of [30], it is proved that $T_{3}$, and $T_{1}$ if $\gamma_{s}>\beta+1/2$, belong to a bounded subset of $H^{\alpha}(\mathbb{R})$ where $\alpha=1/2+\eta$ for some $\eta>0$ small enough. In that paper, they conclude from Proposition 1 in [23] that $B$ is $\mathbb{P}$-Donsker where $\mathbb{P}$ denotes the distribution of the observations they consider. In our case, the underlying distribution $\widehat{\mathbb{P}}_{n}$ depends on $n$ so we cannot use a similar argument. Note however that proving a Donsker property is much stronger than proving tightness, while only tightness is required here. Nevertheless, asymptotic tightness is proven in Lemma 4 below by extenging part of the arguments from [20] to the case of an underlying distribution that depends on $n$. Note that the lemma does not require that the event $E_{n}$ holds.
Lemma 4.
Under the assumptions of Theorem 7, the processes $\sqrt{n}\int T_{j}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ with $j=3$, and $j=1$ if $\gamma_{s}>\beta+1/2$, are asymptotically tight conditionally on the original observations.
To deal with the case $j=2$ we assume that the second event that defines $E_{n}$ in (D.1) is satisfied (see (33) below). Hence, the result still holds if we assume that $E_{n}$ holds for all $n$. In Section 4.2.2 of [30], it is proved that the functions $T_{2}(x)/(1+ix)$ belong to a bounded subset of $H^{\alpha}(\mathbb{R})$ where $\alpha=1/2+\eta$ for some $\eta>0$ small enough. We combine this result with empirical process theory and an entropy bound, taken from [23], for the set $H^{\alpha}(\mathbb{R})$, to obtain the following lemma.
Lemma 5.
If Assumption B holds, then the process $\sqrt{n}\int T_{2}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ is asymptotically tight conditionally on the original observations.
It remains to deal with the case $j=1$ when $\gamma_{s}\leq\beta+1/2$. Fix $\xi>0$ and defined
$K_{h}^{(0)}=K_{h}\mathds{1}_{[-\xi,\xi]}.$
In Section 4.2.3 of [30], it is proved that if $\gamma_{s}>\beta$, and if we define $T$ similarly as $T_{1}$ with $K_{h}-K_{h}^{(0)}$ instead of $K_{h}$, then $T$ is contained in a bounded subset of $H^{\alpha}(\mathbb{R})$ where $\alpha=\gamma_{s}-\beta+1>1/2$. Hence, the same arguments as in the proof of Lemma 4 imply that $\sqrt{n}\int T(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ is asymptotically tight conditionally on the original observations.
It remains to deal with $T_{1}^{(0)}$, that is defined similarly as $T_{1}$ with the truncated kernel $K_{h}^{(0)}$ instead of $K_{h}$. The arguments in Section 4.2.3 of [30] easily generalize to the setting where a fix probability $\mathbb{P}$ is replaced by $\widehat{\mathbb{P}}_{n}$ if we assume that $\hat{f}_{n}(x)\leq C$ for all $x\in\mathbb{R}$. Indeed, our limiting process is the same as that in that paper so we know that it is pre-Gaussian, so in the bootstrap setting it remains to verify the five conditions of Theorem 12 in [24] as in [30], but with $\mathbb{P}$ replaced by $\widehat{\mathbb{P}}_{n}$. Condition (a) does not depend on the underlying probability. To verify condition (b) we write
$Q_{\tau}^{\prime}=\{r-q|r,q\in Q,\left(\int|r-q|^{2}\mbox{d}\widehat{\mathbb{P}}_{n}\right)^{1/2}\leq\tau\}$
and we notice that the bounds in [30] are uniform with respect to the underlying probability $\mathbb{P}$, whence it holds if the underlying probability is $\widehat{\mathbb{P}}_{n}$. Similarly, the check of (c) and (f) relies on bounds that are uniform in the underlying probability. The check of (d) relies on the fact that the density function of the underlying probability is bounded. This condition continues to hold in our setting if we assume that $\hat{f}_{n}(x)\leq C$ for all $x\in\mathbb{R}$. Hence, we conclude that $\sqrt{n}\int T_{1}^{(0)}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ is asymptotically tight conditionally on the original observations. This completes the proof of Theorem 7. $\Box$
D.2 Proof of Lemma 3
We argue conditionally on the original observations. It follows from (25) that for a fixed $t$, $\sqrt{n}S_{nt}$ is the mean of iid random variables $g_{th}(Y_{1}^{*}),\dots,g_{th}(Y_{n}^{*})$ where
$g_{th}(Y_{j}^{*})=\mathcal{F}^{-1}\left[(1/\varphi_{\epsilon})(-\ {\bullet})\mathcal{F}K_{h}\right]\star\zeta_{t}(Y_{j}^{*})$
minus expectation. Note that the variance of those variables is
$v_{nt}$ $=$ $\int g_{th}^{2}\mbox{d}\widehat{\mathbb{P}}_{n}-\left(\int g_{th}\mbox{d}\widehat{\mathbb{P}}_{n}\right)^{2}.$
Since $E_{n}$ holds for all $n$, we have the uniform bounds $\sup_{t}|\hat{f}_{n}(t)|\leq C$, $\|\hat{f}_{n}\star f_{\epsilon}\|_{\infty}\leq C\int f_{\epsilon}(x)\mbox{d}x\leq C$ and
$\int(1+x^{2})^{(2+\delta)/2}\widehat{\mathbb{P}}_{n}(\mbox{d}x)=\int(1+x^{2})^{(2+\delta)/2}\hat{f}_{n}(x)\mbox{d}x\leq C.$
Using these bounds and similar arguments as for the control of the term in Equation (16) of [30], one can prove that there exists $C^{\prime}>0$ such that for all $h$ and $n$,
$\int\ \left|g_{th}(x)\right|^{2+\delta}\widehat{\mathbb{P}}_{n}(\mbox{d}x)\leq C^{\prime}\|\zeta_{t}\|_{Z^{\beta+\delta,1/2+\beta+\delta}}$ (26)
(using the same notations as in [30]), which is finite if $\beta+\delta<1/2$ (see also Example 1 in [30]). Since $\beta<1/2$, on can choose $\delta$ small enough so that the latter condition holds, hence the bound in (26) is finite.
Now, we consider convergence of $v_{nt}$. It follows from the previous display combined to Jensen’s inequality that there exists $C>0$ such that for all integers $n$,
$\int\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x\leq C,$
which means that the functions $\left|g_{th}\right|^{2}\hat{f}_{n}$ all belong to $L_{1}(\mathbb{R})$ with a $L_{1}$-norm bounded by $C$. This implies that for all $\epsilon>0$ one can find $R>0$ independent of $n$ such that
$\sup_{n}\int_{|x|>R}\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x\leq\epsilon,$
see e.g. Theorem 2.32 in [12]. One can choose $R$ in such a way that, in addition,
$\int_{|x|>R}\left|g_{t}(x)\right|^{2}f(x)\mbox{d}x\leq\epsilon,$
where $g_{t}=\mathcal{F}^{-1}\left[(1/\varphi_{\epsilon})(-\ {\bullet})\right]\star\zeta_{t}$, by integrability of $|g_{t}|^{2}f$. Now, the functions $\left|g_{th}\right|^{2}\hat{f}_{n}$ are bounded on [R,R][-R,R]: there exists a constant $A>0$ such that the supremum over $n$ of the sup-norms over [R,R][-R,R] of the functions is bounded above by $A$. The constant $A$ is integrable with respect to the Lebesgue measure on [R,R][-R,R], so it follows from the dominated convergence theorem under the assumptions that $h$ converges to zero and $\hat{f}_{n}$ converges to $f$ that
$\int_{|x|\leq R}\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x-\int_{|x|\leq R}\left|g_{t}(x)\right|^{2}f(x)\mbox{d}x=o(1).$
Hence, splitting the integrals over $\mathbb{R}$ into the sum of integrals over [R,R][-R,R] and $\mathbb{R}\backslash[-R,R]$ yields
$\left|\int\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x-\int\left|g_{t}(x)\right|^{2}f(x)\mbox{d}x\right|$
$\qquad\qquad\leq o(1)+\int_{|x|>R}\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x+\int_{|x|>R}\left|g_{t}(x)\right|^{2}f(x)\mbox{d}x$
$\qquad\qquad\leq o(1)+2\epsilon.$
Letting $n\to\infty$ and then $\epsilon\to 0$ one obtains
$\lim_{n\to\infty}\int\left|g_{th}(x)\right|^{2}\hat{f}_{n}(x)\mbox{d}x=\int\left|g_{t}(x)\right|^{2}f(x)\mbox{d}x.$
Similarly, $\int g_{th}\hat{f}_{n}$ converges to $\int g_{t}f$, so we conclude that $v_{nt}$ converges to $\int\left|g_{t}\right|^{2}f-(\int g_{t}f)^{2},$ which is equal to the variance of the centered Gaussian variable $\mathbb{G}_{F}(t).$ Combining with (26) and the central limit theorem under the Lyapounov condition proves that $\sqrt{n}S_{nt}$ converges in distribution to $\mathbb{G}_{F}(t).$
The proof can be extended (at the price of cumbersome notations and using a vectoriel central limit theorem) to show that for arbitrary fixed $k\in\mathbb{N}$ and $t_{1},\dots,t_{k}$, the joint distribution of $\sqrt{n}S_{nt_{1}},\dots,\sqrt{n}S_{nt_{k}}$ converges to that of $\mathbb{G}_{F}(t_{1}),\dots,\mathbb{G}_{F}(t_{k})$, which yields Lemma 3. $\Box$
D.3 Proof of Lemma 4
Recall that from Section 4.2.2 of [30], we know that $T_{j}$ with $j=3$, and $j=1$ if $\gamma_{s}>\beta+1/2$, belongs to a bounded subset $B$, say, of $H^{\alpha}(\mathbb{R})$ where $\alpha=1/2+\eta$ for some $\eta>0$ small enough. Note that by definition of $T_{j}$, $B$ can be indexed by $t\in\mathbb{R}$:
$B=\{g_{t},\ t\in\mathbb{R}\}.$ (27)
As mentionned in the paragraph before Theorem 1.3 of that paper, $H^{\alpha}$ is a Hilbert space for arbitrary $\alpha\geq 0$ if we fix the seminorm so it is a norm, whence we can consider an orthonormal basis $(f_{k})_{k\in\mathbb{N}}$. Moreover, it follows from Theorem 1.2 of the same paper that for all $\in H^{\alpha}, ff$ is continuous and bounded with
$\|f\|_{\infty}\leq\|u^{-1}\|_{L^{2}}\|f\|_{H^{\alpha}}$ (28)
where $u(x)=\langle x\rangle=(1+x^{2})^{1/2}$. Let $T$ be the inclusion function from $H^{\alpha}$ to the set of continuous and bounded functions on $\mathbb{R}$, whence $T$ is linear with norm that is less than or equal to $\|k^{-1}\|_{L^{2}}$. Let $m\in\mathbb{N}$ to be chosen later. One can obtain Equation (*) on page 325 of [20] with $P$ replaced by the conditionnal probability $\mathbb{P}^{*}$ given $(Y_{j})_{j\in\mathbb{N}}$; and then that
$\sup_{n}\mathbb{P}^{*}\left(n^{-1/2}\sup_{f\in B}\left|\sum_{j=1}^{n}(\delta_{Y_{j}^{*}}-\widehat{\mathbb{P}}_{n})(f)-\Lambda_{m}(\delta_{Y_{j}^{*}}-\widehat{\mathbb{P}}_{n})(f)\right|\geq m^{-1}\right)\leq m^{-1}$ (29)
where for all $g_{t}:=\sum_{k}a_{kt}f_{k}\in B$,
$n^{-1/2}\sum_{j=1}^{n}\Lambda_{m}(\delta_{Y_{j}^{*}}-\widehat{\mathbb{P}}_{n})(g_{t})=\sum_{k=1}^{i(m)}a_{kt}Z_{nk}$
for some sufficiently large $i(m)$, with
$Z_{nk}=n^{-1/2}\sum_{j=1}^{n}(\delta_{Y_{j}^{*}}-\widehat{\mathbb{P}}_{n})(f_{k}).$
Thanks to (27) we can rewrite (29) as
$\sup_{n}\mathbb{P}^{*}\left(\sup_{t\in\mathbb{R}}\left|\mathbb{G}_{n}^{*}(t)-\mathbb{G}_{n}(t)\right|\geq m^{-1}\right)\leq m^{-1}$ (30)
where
$\mathbb{G}_{n}^{*}(t)=n^{-1/2}\sum_{j=1}^{n}(\delta_{Y_{j}^{*}}-\widehat{\mathbb{P}}_{n})(g_{t}),\quad\mathbb{G}_{n}(t)=\sum_{k=1}^{i(m)}a_{kt}Z_{nk}.$ (31)
For arbitrary $k$ one has
$\mathbb{E}^{*}\left(Z_{nk}\right)^{2}$ $\leq$ $\mathbb{E}^{*}\left(f_{k}(Y_{1}^{*})\right)^{2}$
$\leq$ $\|f_{k}\|_{\infty}^{2}$
$\leq$ $\|u^{-1}\|_{L^{2}}^{2}$
using (28) together with the assumption that $f_{k}\in H^{\alpha}$ has norm one. Here $\mathbb{E}^{*}$ denotes the expectation conditionally on $(Y_{j})_{j\in\mathbb{N}}$. This implies that for arbitrary $\epsilon>0$ there exists $C_{\epsilon}>0$ such that
$\sup_{n}\mathbb{P}^{*}\left(|Z_{nk}|>C_{\epsilon}\right)\leq\epsilon.$
Define $K_{k}:=\{a_{k}x,\ |x|\leq C_{\epsilon}\}$ with $x\in\mathbb{R}$, $a_{k}$ the function from $\mathbb{R}$ to $\mathbb{R}$ such that $a_{k}(t)=a_{kt}$. We have $\sup_{t}|a_{k}(t)|\leq\sup_{t}\|g_{t}\|_{H^{\alpha}}<\infty$ since $B$ is bounded whence $a_{k}\in\ell_{\infty}(\mathbb{R})$, and $K_{k}$ is a compact subset of $\ell_{\infty}(\mathbb{R})$ that satisfies
$\mathbb{P}^{*}\left(a_{k}Z_{nk}\in K_{k}\right)\geq 1-\epsilon.$
We conclude that the sequence $a_{k}Z_{nk}$ is asymptotically tight in $\ell_{\infty}(\mathbb{R})$ for arbitrary $k$ and therefore, it follows from [33, Theorem 1.5.7.] that $\mathbb{G}_{n}:=\sum_{k=1}^{i(m)}a_{kt}Z_{nk}$ is asymptotically tight as well in $\ell_{\infty}(\mathbb{R})$.
Now, fix $\epsilon>0$ and let $m$ large enough so that $m^{-1}\leq\epsilon/2$. By asymptotic tightness, we can find a compact $K_{m}$ in the span of $\{a_{k},\ k\leq i(m)\}\subset\ell_{\infty}(\mathbb{R})$ such that for all $\delta>0$,
$\sup_{n}\mathbb{P}^{*}\left(\mathbb{G}_{n}\in K_{m}^{\delta}\right)\geq 1-\epsilon/2$
and in particular,
$\sup_{n}\mathbb{P}^{*}\left(\mathbb{G}_{n}\in K_{m}^{1/m}\right)\geq 1-\epsilon/2.$ (32)
With $\tilde{K}$ a compact subset of $\ell_{\infty}(\mathbb{R})$ that contains $K_{m}^{2/m}$ we then have
[…]
$\leq$ $\mathbb{P}^{*}\left(\|\mathbb{G}_{n}-K_{m}\|_{\infty}>1/m\right)+\mathbb{P}^{*}\left(\|\mathbb{G}_{n}^{*}-\mathbb{G}_{n}\|_{\infty}>1/m\right)$
$\leq$ $\epsilon/2+m^{-1}$
$\leq$ $\epsilon.$
Hence, $\mathbb{G}_{n}^{*}$ is asymptotically tight, where $\mathbb{G}_{n}^{*}=\sqrt{n}\int T_{j}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$. $\Box$
D.4 Proof of Lemma 5
Thanks to Assumption B, we know that for arbitrary $\epsilon>0$, there exists $C_{\epsilon}>0$ such that for all sufficiently large $n$, one has
$\int|x|^{2+\delta}\hat{f}_{n}(x)\mbox{d}x\leq C_{\epsilon}$ (33)
with probability larger than $1-\epsilon$. Hence, it suffices to show that for arbitrary $\epsilon>0$, the process $\sqrt{n}\int T_{2}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})$ is asymptotically tight conditionally on the original observations provided that (33) holds for sufficiently large $n$. For this reason, we assume without loss of generality in the sequel that (33) holds where $\epsilon>0$ is arbitrary.
Recall that from Section 4.2.2 of [30], we know that the functions $T_{2}(x)/(1+ix)$ belong to a bounded subset of $H^{\alpha}(\mathbb{R})$ where $\alpha=1/2+\eta$ for some $\eta>0$ small enough. Hence, the functions $T_{2}$ belong to a bounded subset $B$ say, of the set of functions $f$ such that $\|fu^{-1}\|_{H^{\alpha}}<\infty$ where $u(x)=\langle x\rangle=(1+x^{2})^{1/2}$. In the sequel, we assume $\eta$ small enough so that $\eta<\delta/2$ where $\delta$ is taken from Assumption B, and we consider some $\gamma\in(\eta,\delta/2)$. It follows from (33) that
$\int\langle x\rangle^{2+\delta}\mbox{d}\widehat{\mathbb{P}}_{n}(x)$ $\lesssim$ $1+\int|x|^{2+\delta}\mbox{d}\widehat{\mathbb{P}}_{n}(x)$
$=$ $1+\int|x|^{2+\delta}\hat{f}_{n}(x)\mbox{d}x$
$\leq$ $1+C_{\epsilon}.$
Therefore, it follows from Hölder’s inequality that
$\sup_{n}\int\langle x\rangle^{2+2\gamma}\mbox{d}\widehat{\mathbb{P}}_{n}(x)\leq(1+C_{\epsilon})^{(2+2\gamma)/(2+\delta)}.$
Since $H^{\alpha}=B_{2,2}^{\alpha}$, see e.g. the Appendix in [30], Corollary 1 in [23] (with $p=q=2$, $\beta=-1$, and $s=\alpha$, $r=2$) provides a bound for the bracketing metric entropy of $B$:
$\sup_{n}H_{[]}(\delta,B,\|\ .\ \|_{L_{2}(\widehat{\mathbb{P}}_{n})})\lesssim\delta^{-1/\alpha}.$
Moreover, the third assertion in Proposition 3 of the same paper implies that there exists a real number $K>0$ such that $\sup_{f\in B}|f|\leq Ku$, whence $Ku$ is an envelope of $B$. Due to Hölder’s inequality, the enveloppe satisfies
$\sup_{n}\int(Ku)^{2}\widehat{\mathbb{P}}_{n}(x)$ $=$ $\sup_{n}K^{2}\int\langle x\rangle^{2}\mbox{d}\widehat{\mathbb{P}}_{n}(x)<\infty,$
so Theorem 2.14.1 in [33] yields that if (33) holds for sufficiently large $n$ then
$\sup_{n}\mathbb{E}^{*}\left\|\sqrt{n}\int T_{2}(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})\right\|_{\infty}$ $\lesssim$ $\sup_{n}\mathbb{E}^{*}\sup_{f\in B}\left|\sqrt{n}\int f(\mbox{d}{\mathbb{P}}_{n}^{*}-\mbox{d}\widehat{\mathbb{P}}_{n})\right|$
$\lesssim$ $\int_{0}^{1}\sqrt{1+\epsilon^{-1/\alpha}}\mbox{d}\epsilon$
$\lesssim$ $1+\int_{0}^{1}\delta^{-1/(2\alpha)}\mbox{d}\delta$
which is finite since $\alpha>1/2$. The lemma follows. $\Box$
Appendix E Proofs for Section 5.1
E.1 Proof of Lemma 2
It follows from the second claim of Lemma 1 that $\hat{F}_{n}^{d}$ is well defined with $\hat{F}_{n}^{d}(t)=0$ for all $t\leq 0$ and
$\lim_{t\to\infty}\hat{F}_{n}^{d}(t)=1$
by definition. Moreover, $\hat{F}_{n}^{d}$ is continuous by concavity and non-decreasing by Theorem 1, whence the lemma follows. $\Box$
E.2 Proof of Theorem 6
Recall that from Example 1 in [30], we know that we can choose $\zeta=\mathds{1}_{(-\infty,0]}$ in Theorem 7 provided that $\gamma_{s}<1/2$ and in that case, $\vartheta_{t}^{*}$ is the common distribution function of the iid observations while $\widehat{\vartheta}_{t}^{*}$ is the corresponding deconvolution estimator based on the bootstrap sample. Recall moreover that thanks to Corollary 1 we know that $\hat{f}_{n}$ uniformly converges to $f_{0}$. In order to prove Theorem 6, we will need to apply Theorem 7 with the above choice of $\zeta$, and $f$ replaced by $f_{0}$ in Assumption B, and also Lemma 2 so the first step is to show that the assumptions of theorems 2 and 7 are satisfied under the assumptions of Theorem 6. In the following Lemma, which proof is given in Section E.3, we show that assumptions of Theorem 2 hold.
Lemma 6.
  1. 1.
    Assumptions D’ and E’ imply Assumptions D and E.
  2. 2.
    Let $\gamma$ and $\beta$ be positive numbers such that $\gamma>\beta$, and $h_{n}$ a sequence of positive numbers that converges to zero as $n\to\infty$. If $h_{n}^{2+2\beta}n\to\infty$ then $h_{n}^{\rho}n\to\infty$ for some $\rho>4\beta-4\gamma+2$.
  3. 3.
    The kernel $K$ given by (13) satisfies conditions K for all $L<rs$.
In the sequel we denote by $\hat{f}_{n}$ the slope of the distribution function $\mathcal{M}\hat{F}_{n}^{d}$ defined by (6) and by $f_{n}$ the slope of $\mathcal{M}F_{n}^{d}$, which are well defined almost everywhere by concavity. Note that thanks to Lemma 2, $\hat{f}_{n}$ is a genuine density function (with probability that tends to one) in the sense that it is non-negative and integrates to one, which is not necessarily the case of $f_{n}$. Since we are concerned with convergence in probablity, we can restrict ourselves to a sequence of events whose probability tends to one as $n\to\infty$ so we assume in the sequel without loss of generality that $\hat{f}_{n}$ is a genuine density function. Hence, it follows from Theorem 7 (where we set $\zeta=\mathds{1}_{(-\infty,0]}$ so that $\vartheta_{t}^{*}=\mathcal{M}\hat{F}_{n}^{d}(t)$ and $\hat{\vartheta}_{t}^{*}=\widetilde{F}_{m}^{d*}(t)$, see also Example 1 in [30]) that the first claim of Theorem 6 holds provided that Assumption B is satisfied with $f$ replaced by $f_{0}$. This means that the first claim is a direct consequence of Corollary 1, Lemma 6 and the following lemma, that is proved in Section E.4.
Lemma 7.
Under the assumptions of Theorem 6, with $\hat{f}_{n}$ the slope of the distribution function $\mathcal{M}\hat{F}_{n}^{d}$ defined by (6), one has (20), (21) and (22).
We turn to the proof of the second claim in Theorem 6, hence we assume that $F$ is concave. As in Section B we consider the operator $\mathcal{W}=\mathcal{M}$-$\mathcal{I}$ where $\mathcal{I}$ stands for the identity operator. Since $\hat{F}_{n}^{d}$ is concave on $\mathbb{R}^{+}$, we have $\mathcal{M}\hat{F}_{n}^{d}=\hat{F}_{n}^{d}$, so the bootstrap version $T_{m}(F_{m}^{d*})$ of the test statistic becomes
$T_{m}(F_{m}^{d*})=\sqrt{m}\|\mathcal{W}F_{m}^{d*}-\mathcal{W}\hat{F}_{n}^{d}\|_{\infty}.$
It follows from Theorem 2 combined with the continuous mapping theorem that
$\sqrt{n}\sup_{t\geq 0}|F_{n}^{d}(t)-F(t)|\stackrel{{\scriptstyle L}}{{\longrightarrow}}\sup_{t\geq 0}|\mathbb{G}_{F}(t)-\mathbb{G}_{F}(0)|$
as $n\to\infty$ with $\mathbb{G}_{F}$ a centered Gaussian Borel random variable in $\ell^{\infty}(\mathbb{R})$, whence the left-hand side is bounded in probability. Hence, with $(\epsilon_{n})_{n}$ an arbitrary sequence of positive numbers that converges to zero as $n\to\infty$ with $\lim_{n\to\infty}\sqrt{n}\epsilon_{n}=+\infty$, one has
$\|F_{n}^{d}-F\|_{\infty}\leq\epsilon_{n}$ (34)
with probability that tends to one. Moreover,
$\left|\lim_{t\to\infty}F_{n}^{d}(t)-1\right|$ $=$ $\left|\lim_{t\to\infty}F_{n}^{d}(t)-\lim_{t\to\infty}F(t)\right|$
$\leq$ $\|F_{n}^{d}-F\|_{\infty}$
whence we have
$\left|\lim_{t\to\infty}F_{n}^{d}(t)-1\right|$ $\leq$ $\epsilon_{n}$ (35)
with probability that tends to one. Combining with the fact that the test statistic in (18) is bounded in probability where $\mathcal{W}F=0$, one gets
$\sqrt{m}\|\mathcal{W}F_{n}^{d}-\mathcal{W}\hat{F}_{n}^{d}\|_{\infty}=O_{\mathbb{P}}(m^{1/2}n^{-1/2}\epsilon_{n}).$}
Since $m\ll n$ one then gets
$T_{m}(F_{m}^{d*})=\sqrt{m}\|\mathcal{W}F_{m}^{d*}-\mathcal{W}F_{n}^{d}\|_{\infty}+o_{\mathbb{P}}(1)$
$=\sqrt{m}\|\mathcal{W}F_{m}^{d*}-\mathcal{W}F\|_{\infty}+o_{\mathbb{P}}(1),$
using again that the test statistic in (18) is bounded in probability for the last equality. Since for arbitrary functions $\theta\in\ell_{\infty}(\mathbb{R}^{+})$ and $a\in\mathbb{R}$ one has $\mathcal{M}(\theta+a)=\mathcal{M}\theta+a$, which implies that $\mathcal{W}(\theta+a)=\mathcal{W}\theta$, we also have
$T_{m}(F_{m}^{d*})=\sqrt{m}\|\mathcal{W}\widetilde{F}_{m}^{d*}-\mathcal{W}F\|_{\infty}+o_{\mathbb{P}}(1).$
In fact, the assumption that that $m\ll n$ is made precisely with the aim to replacing $F_{n}^{d}$ by $F$ in the above displays, and also in (14) which, thanks to (8) and Marshall’s inequality, can be reformulated as
$\sqrt{m}(\widetilde{F}_{m}^{d*}-F)\rightsquigarrow\mathbb{G}_{F}\ \ in\ \ell_{\infty}(\mathbb{R}),$
as $n\to\infty$. Now, we have seen in Section B that $\mathcal{W}$ is Hadamard directionally differentiable at $F$ with derivative $\mathcal{W}_{F}^{\prime}=\mathcal{M}_{F}^{\prime}$-$\mathcal{I}$ so one gets
$\sqrt{m}(\mathcal{W}\widetilde{F}_{m}^{d*}-\mathcal{W}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}\ \ in\ \ell_{\infty}(\mathbb{R}),$
as $n\to\infty$, conditionally. Since the sup-norm is a continuous function on $\ell_{\infty}(\mathbb{R})$, invoking the continuous mapping theorem completes the proof of the second claim in Theorem 6.
It remains to prove the third claim. It follows from the triangle inequality that
$\|\mathcal{M}F_{m}^{d*}-F_{m}^{d*}\|_{\infty}$ $\leq$ $\|\mathcal{M}F_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d}\|_{\infty}+\|\mathcal{M}\hat{F}_{n}^{d}-F_{m}^{d*}\|_{\infty}$
where $\mathcal{M}\hat{F}_{n}^{d}=\mathcal{M}\mathcal{M}\hat{F}_{n}^{d}$ by concavity. Marshall’s inequality [14, Exercise 3.1] then yields
$\|\mathcal{M}F_{m}^{d*}-F_{m}^{d*}\|_{\infty}$ $\leq$ $\|F_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d}\|_{\infty}+\|\mathcal{M}\hat{F}_{n}^{d}-F_{m}^{d*}\|_{\infty}$
$=$ $2\|F_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d}\|_{\infty}$
$\leq$ $2\|\widetilde{F}_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d}\|_{\infty}+2|\widetilde{F}_{m}^{d*}(0)|$
Now, it follows from the continuous mapping theorem combined to Theorem 6 that $\sqrt{m}\|\widetilde{F}_{m}^{d*}-\mathcal{M}\hat{F}_{n}^{d}\|_{\infty}$ converges in distribution to $\|\mathbb{G}_{F_{0}}\|_{\infty}$ while $\sqrt{m}|\widetilde{F}_{m}^{d*}(0)|$ converges in distribution to $|\mathbb{G}_{F_{0}}(0)|$, conditionally as $n\to\infty$. This implies that conditionally, the variable in the last line of the previous display is bounded in probability. This implies that $\sqrt{m}\|\mathcal{M}F_{m}^{d*}-F_{m}^{d*}\|_{\infty}$ is dominated by a variable that is bounded in probability, hence it is bounded in probability as well, conditionally. This completes the proof of Theorem 6. $\Box$
E.3 Proof of Lemma 6
Suppose that Assumptions D’ and E’ hold. Then there exists a positive $C$ such that $\mathbb{E}(X^{8})\leq C$ and $\mathbb{E}(\varepsilon^{8})\leq C$. By Jensen’s inequality, this implies that $\int\left|x\right|^{2+\delta}f\left(x\right)dx<\infty$ and $\int\left|x\right|^{2+\delta}f_{\varepsilon}\left(x\right)dx<\infty$ for all positive $\delta\leq 6.$ Hence, E and D hold.
The second claim follows from the fact that $2+2\beta>4\beta-4\gamma+2$ for all $\gamma>\beta$.
We turn to the third claim. By construction, $\mathcal{F}K$ is compactly supported in [1,1][-1,1], which implies that $K$ is band-limited. Since $\mathcal{F}K$ is even, $K$ is real-valued and symmetric. Moreover, as a compactly supported polynomial, $\mathcal{F}K\in C_{c}^{\infty}(\mathbb{R})$, the space of smooth functions with compact support, and hence its inverse Fourier transform $K$ belongs to the Schwartz space $\mathcal{S}(\mathbb{R})$, defined as the space of all functions that are infinitly differentiable and such that for all non-negative integers $m$ and $n$, the $n$th derivative $f^{(n)}$ is such that the function $x\mapsto|x^{m}f^{(n)}(x)|$ is bounded. This implies that $K\in L^{1}(\mathbb{R})\cap\ell_{\infty}(\mathbb{R})$ whence the first part of assumption K is satisfied. Now, we use the identity
$\int x^{j}K(x)\,dx=(-i)^{j}\left.\frac{d^{j}}{dt^{j}}\mathcal{F}K(t)\right|_{t=0},$
which holds since $\mathcal{F}K\in C_{c}^{\infty}(\mathbb{R})$. By construction, the right-hand side is equal to zero for all $j=1,\dots,rs-1$, and thus $\int x^{j}K(x)\,dx=0$ for $j=1,\dots,L$, for any integer $L<rs$. Since $K\in\mathcal{S}(\mathbb{R})$, all the absolute moments of $K$ are finite so the second part of assumption K holds, as well as the third part, using that $\int K=\mathcal{F}K(0)=1$. $\Box$
E.4 Proof of Lemma 7
In the sequel we denote by $f_{n}^{d}$ the slope of the deconvolution estimator $F_{n}^{d}$ and by $\hat{f}_{n}^{d}$ the slope of $\hat{F}_{n}^{d}$ defined by (4), whence
$\hat{f}_{n}^{d}(t)=\frac{f_{n}^{d}(t)}{\lim_{t\to\infty}F_{n}^{d}(t)}$ (36)
for all $t$. We prove in Sections E.4.1 and E.4.2 below that under the assumptions of Theorem 6 one has
$\sup_{t\in\mathbb{R}^{+}}|f_{n}^{d}(t)|=O_{p}(1),$ (37)
and because $r>4$ there exists $\delta>0$ such that
$\int_{0}^{\infty}|t|^{2+\delta}f_{n}^{d}(t)dt=O_{p}(1).$ (38)
Using those properties of the estimator, the equations (20), (21) and (22) are proved in the subsequent sections E.4.3, E.4.4 and E.4.5, respectively. $\Box$
E.4.1 Proof of (37)
Recall that by assumption, $f\in H^{\alpha}\left(\mathbb{R}\right)$ for some $\alpha>1/2$. It follows from the Cauchy-Schwarz’s inequality that
$\int|\mathcal{F}f|$ $=$ $\int(1+x^{2})^{\alpha/2}|\mathcal{F}f(x)|\times(1+x^{2})^{-\alpha/2}dx$
$\leq$ $\left(\int(1+x^{2})^{\alpha}|\mathcal{F}f(x)|^{2}dx\right)^{1/2}\times\left(\int(1+x^{2})^{-\alpha}dx\right)^{1/2},$
where the first term on the last line is bounded because $f\in H^{\alpha}\left(\mathbb{R}\right)$, and the second one if finite since $\alpha>1/2$. Hence,
$\int|\mathcal{F}f|<\infty$ (39)
is finite so we can apply Corollary 1 to obtain that $\|f_{n}^{d}-f\|_{\infty}=o_{p}(1)$. Combining with boundedness of $f$ and the triangle inequality proves that $\|f_{n}^{d}\|_{\infty}=O_{p}(1)$. $\Box$
E.4.2 Proof of (38)
Let $\delta>0$ be sufficently small so that $2+\delta<4$. Let $q=2+\delta$, $k=1$, $v_{q}=\mathbb{E}(|X|^{q})$ and $\hat{v}_{q}=\int_{-\infty}^{\infty}|t|^{q}\,f_{n}^{d}(t)dt$. Under assumption (5) one has
$\left|\left(1/\varphi_{\varepsilon}\right)(t)\right|\leq C^{\prime}(1+|t|^{2})^{\beta/2},\quad\text{for all }t\in\mathbb{R}.$
Hence, under our Assumptions D’ and E’ we have $F_{\epsilon}\in\mathcal{F}_{5}(\beta,C_{1})$ and $F_{X}\in\mathcal{G}_{6}(0,C_{2})$ in the notations of [15]. Moreover, the assumption that $r>4$ implies that $r>\max(q,4)$ so it follows from Theorem 3.6 in [15] that
$\mathbb{E}[(\hat{v}_{q}-v_{q})^{2}]\leq C\left(h^{2q}+n^{-1}h^{-(2\beta-2q-1)}\right).$
The upper bound tends to zero since $h\to 0$ and $\beta<1/2$, so it follows from Markov’s inequality that $(\hat{v}_{q}-v_{q})^{2}=o_{P}(1)$, whence $\hat{v}_{q}=v_{q}+o_{P}(1)$. From the assumption that $v_{8}$ is finite, we can deduce that $v_{q}$ is finite whence $\hat{v}_{q}=O_{P}(1)$, which proves (38). $\Box$
E.4.3 Proof of (20)
Since $\hat{f}_{n}$ is the slope of the least concave majorant of $\hat{F}_{n}^{d}$ and $\hat{F}_{n}^{d}(0)=0$ by definition, we have
$\hat{f}_{n}(0)$ $=$ $\sup_{t>0}\frac{\hat{F}_{n}^{d}(t)}{t}.$
Moreover, it follows from Lemma 2 that $\hat{f}_{n}$ is a genuine density function, in particular, it is non-negative at every point (with probability that tends to one). Since (35) holds with probablity that tends to one, it follows from the definition (4) that
$0\leq\hat{f}_{n}(0)$ $\leq$ $\sup_{t>0}\frac{F_{n}^{d}(t)}{t}\times\frac{1}{1-\epsilon_{n}}$ (40)
with probability that tends to one. Here, $(\epsilon_{n})_{n}$ is an arbitrary sequence of positive numbers that converges to zero as $n\to\infty$ with $\lim_{n\to\infty}\sqrt{n}\epsilon_{n}=+\infty$. We will split the above supremum into two parts. We begin with the supremum near zero. Let $T>0$. Since $f_{n}^{d}$ is the derivative of $F_{n}^{d}$ we have
$\sup_{t\in(0,T]}\frac{F_{n}^{d}(t)}{t}$ $=$ $\sup_{t\in(0,T]}\left(\frac{1}{t}\int_{0}^{t}f_{n}^{d}(x)dx\right)$
$\leq$ $\sup_{t\in(0,T]}|f_{n}^{d}(t)|=O_{p}(1),$
where we used (37) for the last equality. We now consider the supremum far from zero. Since (34) holds with probability that tends to one we obtain
$\sup_{t\geq T}\frac{F_{n}^{d}(t)}{t}$ $\leq$ $\frac{1}{T}\left(F(t)+\epsilon_{n}\right)$
with probablity that tends to one, so that the supremum on the left hand side is bounded in probability. The supremum over all $t>0$ is bounded from above by the maximum between the supremum near zero and the supremum far from zero so we obtain that
$\sup_{t>0}\frac{F_{n}^{d}(t)}{t}=O_{p}(1).$
Combining with (40) yields that $\hat{f}_{n}(0)=O_{p}(1).$ Now $\hat{f}_{n}$ is a non-increasing function on $0,\infty)$ as being the slope of a concave function $\mathcal{M}F_{n}^{d}$ whence,
$\sup_{x}\hat{f}_{n}(x)\leq\hat{f}_{n}(0)=O_{p}(1).$
This completes the proof of (20). $\Box$
E.4.4 Proof of (21)
The following proof is inspired by Section 1.A.1 in [27]. Because the function $g:t\mapsto|t|^{2+\delta}$ is non-decreasing and takes value zero at point zero, it is possible, for each $k\in\mathbb{N}$, to define a sequence of upper sets (that is, of open or closed right half lines) $U_{1},\dots,U_{k}$, and a sequence of positive numbers $a_{1},\dots,a_{k}$ (all of which may depend on $k$) such that as $k\to\infty$ one has
$g(t)=\lim_{k\to\infty}\sum_{i=1}^{k}a_{i}\mathds{1}_{U_{i}}(t)$
for all $t$, and the approximating functions $g_{k}:=\sum_{i=1}^{k}a_{i}\mathds{1}_{U_{i}}$ are such that $g_{k}\geq g_{\ell}$ for all $k\geq\ell$. To alleviate the notation we write $M_{n}=\mathcal{M}\hat{F}_{n}^{d}$, so that $\hat{f}_{n}$ is the slope of $M_{n}$. Recall that from Lemma 2, we can assume without loss of generality that $\hat{f}_{n}$ is a probability density function, whence it is non negative and integrates to one. It follows from the monotone convergence theorem that
$\int|t|^{2+\delta}\hat{f}_{n}(t)dt$ $=$ $\lim_{k\to\infty}\int\sum_{i=1}^{k}a_{i}\mathds{1}_{U_{i}}(t)\hat{f}_{n}(t)dt$
$=$ $\lim_{k\to\infty}\sum_{i=1}^{k}a_{i}(1-M_{n}(u_{i}))$
where $u_{i}$ denotes the left hand boundary of $U_{i}$. But $M_{n}$ is above $\hat{F}_{n}^{d}$ as being the least concave majorant of the latter function so the previous equalities yield
$\int|t|^{2+\delta}\hat{f}_{n}(t)dt$ $\leq$ $\lim_{k\to\infty}\sum_{i=1}^{k}a_{i}(1-\hat{F}_{n}^{d}(u_{i})).$
Now, the slope of $\hat{F}_{n}^{d}$ is $\hat{f}_{n}^{d}$ that integrates to one, so $(1-\hat{F}_{n}^{d}(u_{i}))=\int\mathds{1}_{U_{i}}(t)\hat{f}_{n}^{d}(t)dt$ which yields
$\int|t|^{2+\delta}\hat{f}_{n}(t)dt$ $\leq$ $\lim_{m\to\infty}\int\sum_{i=1}^{m}a_{i}\mathds{1}_{U_{i}}(t)\hat{f}_{n}^{d}(t)dt$
$=$ $\int g(t)\hat{f}_{n}^{d}(t)dt,$
using again the monotone convergence theorem (possibly decomposing $\hat{f}_{n}^{d}$ into the sum of its positive and negative parts) for the last equality. Combining with (36) and (34), that holds with probability that tends to one for some $\epsilon_{n}\to 0$, then yields
$\int|t|^{2+\delta}\hat{f}_{n}(t)dt$ $\leq$ $\frac{1}{1-\epsilon_{n}}\left|\int g(t)f_{n}^{d}(t)dt\right|$
with probability that tends to one. Combining with (38) completes the proof of (21). $\Box$
E.4.5 Proof of (22)
Let $\alpha\in(0,1/2)$. Since $\hat{f}_{n}$ is monotone, it follows from integration by parts that
$\mathcal{F}\hat{f}_{n}(x)$ $=$ $\left[\frac{e^{iux}}{ix}\hat{f}_{n}(u)\right]_{0}^{\infty}-\frac{1}{ix}\int_{0}^{\infty}e^{iux}d\hat{f}_{n}(u).$
Moreover, Lemma 2 ensures that $\hat{f}_{n}$ is non negative and integrates to one with probability that tends to one so (because one can restrict without loss of generality to an event whose probability tends to one in order to prove boundedness in probability) we can assume without loss of generality that $\lim_{u\to\infty}\hat{f}_{n}(u)=0$, whence
$|\mathcal{F}\hat{f}_{n}(x)|$ $=$ $\left|-\frac{1}{ix}\hat{f}_{n}(0)-\frac{1}{ix}\int_{0}^{\infty}e^{iux}d\hat{f}_{n}(u)\right|$
$\leq$ $\frac{1}{|x|}|\hat{f}_{n}(0)|+\frac{1}{|x|}\int_{0}^{\infty}d\hat{f}_{n}(u),$
using that $|e^{iux}|=1$. Hence,
$|\mathcal{F}\hat{f}_{n}(x)|$ $\leq$ $\frac{1}{|x|}V_{n}.$
where $V_{n}:=|\hat{f}_{n}(0)|+1$ can be assumed to be finite thanks to (20). Since we also have $|\mathcal{F}\hat{f}_{n}(x)|\leq 1$ for all $x$ we conclude that
$\int\langle u\rangle^{2\alpha}|\mathcal{F}\hat{f}_{n}(u)|^{2}du$ $\leq$ $2+\int_{|u|>1}\langle u\rangle^{2\alpha}|\mathcal{F}\hat{f}_{n}(u)|^{2}du$
$\leq$ $2+V_{n}^{2}\int_{|u|>1}\langle u\rangle^{2\alpha}u^{-2}du.$
The integral on the last line is finite since $2\alpha-2<-1$, which completes the proof that (22) holds for all $\alpha\in(0,1/2)$. $\Box$